### v2

# references.bib

Abstract.
Using the projection complex machinery, Bestvina-Bromberg-Fujiwara, Hagen-Petyt and Han-Nguyen-Yang prove that several classes of nonpositively-curved groups admit equivariant quasi-isometric embeddings into finite products of quasi-trees, i.e. having property (QT). In this paper, we unify and generalize the above results by establishing a sufficient condition for relatively hierarchically hyperbolic groups to have property (QT).
As applications, we show that a group has property (QT) if it is residually finite and belongs to one of the following classes of groups: admissible groups, hyperbolic–$2$–decomposable groups with no distorted elements, Artin groups of large and hyperbolic type. We also introduce a slightly stronger version of property (QT), called property (QT’), and show the invariance of property (QT’) under graph products.
Key words and phrases:
Hierarchically hyperbolic, quasi-tree, projection complex, residually finite.
1991 Mathematics Subject Classification:
20F65, 20F67
1. Introduction
Group actions on quasi-trees have been studied intensively in recent years. A quasi-tree is a connected graph with a path metric quasi-isometric to a simplicial tree. In particular, a quasi-tree has asymptotic dimension at most $1$. We say that a finitely generated group $G$ has property (QT) if $G$ equivariantly quasi-isometrically embeds in a finite product of quasi-trees $\prod_{i=1}^{n}T_{i}$ equipped with $\ell^{1}$–metric. Here, “equivariantly” means that $G$ acts on $\prod_{i=1}^{n}T_{i}$ such that the embedding is an orbit map. Such an embedding is called a (QT) embedding of $G$. It is clear that property (QT) is a stronger form of finite asymptotic dimension. Examples of groups with property (QT) include
  • Coxeter groups [DJ99];
  • Residually finite hyperbolic groups [BBF21];
  • Mapping class groups[BBF21];
  • Virtually colorable hierarchically hyperbolic groups whose associated hyperbolic spaces are all quasi-trees [HP22] (including virtually compact special groups [BHS17b], the genus $2$ handlebody group [Che22], fundamental groups of non-geometric graph manifolds [HRSS24], etc.);
  • Fundamental groups of compact orientable $3$–manifolds whose sphere-disk decomposition does not support either Sol or Nil geometry [HNY21];
along with their undistorted subgroups.
The last four examples in the above list are proved to have property (QT) with the help of the projection complex techniques developed in [BBF15, BBFS19]. In particular, property (QT) of mapping class groups strengthens [BBF15, Theorem C], which says that mapping class groups equivariantly quasi-isometrically embeds in a finite product of hyperbolic graphs (of finite asymptotic dimension). Counterexamples of property (QT) include certain special linear groups [Man06, Man08], generalized Baumslag-Solitar groups with infinite monodromy [But22] and groups with Property hereditary (NL) [BFG24]. For some basic corollaries of property (QT), see [HNY21, §2.1&2.2]. Recently, Vergara [Ver24] proves that any finitely generated group with property (QT) has a proper uniformly Lipschitz affine action on $\ell^{1}$ with quasi-isometrically embedded orbits.
In application, we also consider a slightly stronger property, which we call property (QT’). A group $G$ has property (QT’) if $G$ has property (QT) with respect to a finite product of quasi-trees $\prod_{i=1}^{n}T_{i}$ and the $G$–action on $\prod_{i=1}^{n}T_{i}$ is diagonal. By [HNY21, Theorem 1.5], if a residually finite group $G$ is relatively hyperbolic such that every peripheral subgroup has property (QT’), then $G$ has property (QT). Without ambiguity, we also say a $G$–action on a metric space $X$ has property (QT’) if $G$ acts diagonally on a finite product of quasi-trees $\prod_{i=1}^{n}T_{i}$ and $X$ admits a $G$-equivariant quasi-isometric embedding into $\prod_{i=1}^{n}T_{i}$. In particular, if $X$ itself is a finite product of quasi-trees, then any diagonal action on $X$ has property (QT’).
As a generalization of the Masur–Minsky machinery [MM99, MM00], hierarchically hyperbolic groups [BHS17b, BHS19], abbreviated as HHGs, have become an important bridge between mapping class groups, cubical groups, and many other nonpositively-curved groups. A list of papers in this field can be found in [HRSS24]. Coarsely speaking, an HHG is a finitely generated group $G$ whose geometry can be recovered from $G$–equivariant projections to a specified (infinite) collection of hyperbolic spaces. As shown in [BHS17a], HHGs have finite asymptotic dimension. We ask
Question 1.1.
Does every (virtually colorable) HHG have property (QT)?
In this paper, we give a sufficient condition for relative HHGs to have property (QT). Our result unifies and generalizes those of [BBF21], [HP22], and [HNY21] and produces new examples with property (QT). Furthermore, we provide a sufficient condition for the existence of a quasi-median (QT) embedding in the sense of [HP22]. Since this paper first appeared, the existence of a quasi-median (QT) embedding has been used to prove the existence of globally stable cylinders by Petyt, Spriano, and Zalloum [PSZ25]. For background on relative HHGs, we refer the reader to Section 2.3.
Definition 1.2.
Let $(G,\mathfrak{S})$ be a relative HHG. For any $U\in\mathfrak{S}$, We write $G_{U}<\operatorname{Aut}(\mathfrak{S}_{U})$ to mean the image of $\operatorname{Stab}_{G}(U)$ under the restriction homomorphism.
  1. (1)
    We say a domain $U\in\mathfrak{S}$ is of Type I if the following holds.
    1. (a)
      (Hyperbolicity) $\mathcal{C}U$ is hyperbolic.
    2. (b)
      (Acylindrical image) $G_{U}$ acts on $\mathcal{C}U$ acylindrically.
    3. (c)
      (Cobounded nested region) $G_{U}$ acts on $\mathbf{F}_{U}$ coboundedly.
    4. (d)
      (Separable quasi-axes) For any element $g\in\operatorname{Stab}_{G}(U)$ that acts loxodromically on $\mathcal{C}U$, the elementary closure $EC(g)$ is separable in $G$, i.e. $EC(g)$ equals the intersection of all finite-index subgroups of $G$ that contain $EC(g)$.
  2. (2)
    We say a domain $U\in\mathfrak{S}$ is of Type II if the action $G_{U}\curvearrowright\mathcal{C}U$ has property (QT’).
For any $U\in\mathfrak{S}$ of Type II, property (QT’) provides quasi-trees $T_{U}^{i}$ along with $G_{U}$–equivariant maps $\iota^{i}_{U}:\mathcal{C}U\to T_{U}^{i}$ for $i=1,\dots,n_{U}$ such that
$\prod_{i=1}^{n_{U}}\iota_{U}^{i}:\mathcal{C}U\to\prod_{i=1}^{n_{U}}T_{U}^{i}$
is a quasi-isometric embedding.
Our main theorem is as follows.
Theorem 1.3.
Let $(G,\mathfrak{S})$ be a relative HHG that is virtually colorable. If every $U\in\mathfrak{S}$ is of Type I or Type II, then $G$ has property (QT).
Moreover, if for any $D\geq 1$, there exists $D^{\prime}\geq 1$ such that for every $U\in\mathfrak{S}$ of Type II and for each $i=1,\dots,n_{U}$, the map $\iota^{i}_{U}:\mathcal{C}U\to T_{U}^{i}$ sends $(D,D)$–quasi-geodesics to unparametrized $(D^{\prime},D^{\prime})$ quasi-geodesics, then $G$ is coarse median and the (QT) embedding of $G$ is quasi-median.
Sketch of proof.
We roughly explain how to prove Theorem 1.3 in the case that $(G,\mathfrak{S})$ is an HHG with only Type I domains excluding the “moreover” part. This case contains most of the key ideas.
First, we introduce a class of “thick” distances on $G$ each of which is defined using a class of “thick” segments of hierarchy paths on $G$. We prove in Section 3 a thick distance formula saying that the word metric of $G$ can be recovered by summing up these thick distances. This is an analogue of the distance formula for HHGs.
Then we show that any class of thick segments is cofinite up to the group action. Furthermore, these thick segments can be extended to a cofinite collection of quasi-axes. Using projections to these quasi-axes, we can estimate the thick distance. This step is done in Section 4
Finally, we take a finite-index subgroup of $G$, say $H$, such that a collection of quasi-axes as above is divided into finitely many $H$–orbits. Each $H$–orbit satisfies the Bestvina-Bromberg-Fujiwara projection axioms so it gives us a quasi-tree with an $H$–action. We prove that $H$ equivariantly quasi-isometrically embeds in the product of these finitely many quasi-trees in Section 5. Since property (QT) is commensurably invariant, $G$ has property (QT). ∎
As stated in the above proof, property (QT) is commensurably invariant [BBF21, §2.2]. It follows that the conclusion of Theorem 1.3 also holds for any group that is virtually a relative HHG that satisfies our condition, even though such a group may not be a relative HHG itself [PS23].
For most examples of HHGs that emerged from the study, every domain satisfies the first three conditions of Type I. Virtual colorability is also common in practice. Therefore, the biggest restriction of our theorem comes from the assumption of separable quasi-axes. We further discuss it in Section 6 and give an easy-to-use criterion for having separable quasi-axes.
It is clear that residually finite hyperbolic groups and virtually colorable hierarchically hyperbolic groups whose associated hyperbolic spaces are all quasi-trees satisfy the assumption of Theorem 1.3. The following theorem is a collection of more applications in Section 7. Note that these results are new except for mapping class groups.
Theorem 1.4.
The following groups have property (QT).
  • Mapping class groups;
  • Residually finite admissible graphs of groups;
  • Residually finite hyperbolic–$2$–decomposable groups with no distorted elements;
  • Residually finite Artin groups of large and hyperbolic type;
  • Graph products of groups with every vertex group having property (QT’).
Moreover, the (QT) embeddings for the first four classes of groups are quasi-median.
Acknowledgements.
The author is grateful to his PhD supervisor, Koji Fujiwara, for many helpful discussions on this paper. He thanks Wenyuan Yang for discussing with him the paper [HNY21] that became the motivation for this work. He thanks Shengkui Ye for his comments on residually finite Artin groups. This work was supported by JST SPRING, Grant Number JPMJSP2110.
2. Background
2.1. Quasi-isometric embeddings and acylindricity
Given constants $\lambda\geq 1$, $c\geq 0$, we say that a (coarse) map $f:X\to Y$ between metric spaces $(X,\operatorname{\mathsf{d}}_{X})$ and $(Y,\operatorname{\mathsf{d}}_{Y})$ is a $(\lambda,c)$quasi-isometric embedding if
$\frac{1}{\lambda}\operatorname{\mathsf{d}}_{X}(x_{1},x_{2})-c\leq\operatorname{\mathsf{d}}_{Y}(f(x_{1}),f(x_{2}))\leq\lambda\operatorname{\mathsf{d}}_{X}(x_{1},x_{2})+c$
for all $x_{1},x_{2}\in X$. A $(\lambda,c)$–quasi-isometric embedding $\gamma:[0,l]\to X$ is called a (parametrized) $(\lambda,c)$quasi-geodesic in $X$. A coarse map $\gamma:[0,l]\to X$ is an unparametrized $(\lambda,c)$–quasi-geodesic if there is a strictly increasing function $f:[0,l]\to[0,l]$ with $f(0)=0$, $f(l)=l$ such that $\gamma\circ f$ is a $(\lambda,c)$–quasi-geodesic. We also use the term “quasi-geodesic” to mean a quasi-isometric embedding of $\mathbb{R}$. We will not distinguish between a quasi-geodesic and its image in $X$.
A geodesic metric space is called $\delta$hyperbolic (or simply, hyperbolic) for $\delta\geq 0$ if for any geodesics $\alpha,\beta,\gamma$ that form a triangle, $\alpha$ is contained in the $\delta$–neighborhood of $\beta\cup\gamma$ [Gro87]. For a $\delta$–hyperbolic space $X$, an isometry $g:X\to X$ is called loxodromic if the $g$–orbit $n\mapsto g^{n}x$ is a quasi-geodesic for some (equivalently, for any) $x\in X$.
Let $X$ be a hyperbolic space and $G$ be a group acting by isometries on $X$ with a loxodromic element $g$. Given constants $\lambda\geq 1$, $c\geq 0$, a $(\lambda,c)$–quasi-geodesic $\gamma\subset X$ is called a $(\lambda,c)$quasi-axis for $g$ if $\gamma$ is $g$–invariant. The elementary closure of $g$ in $G$, $EC_{G}(g)$, is the subgroup of $G$ that stabilizes $\gamma$ up to bounded Hausdorff distance. If there is no ambiguity in $G$, we often simplify the notation as $EC(g)$. Equivalently, it is the stabilizer of the set $\gamma(\pm\infty)$, the points at infinity of $\gamma$. Thus, the elementary closure does not depend on the choice of $\gamma$. Everything that commutes with $g$ is contained in $EC(g)$ (including powers and roots), but there may be other elements.
A group action $G\curvearrowright X$ by isometries is called acylindrical [Bow08] if for any $r\geq 0$, there exist constants $R,N\geq 0$ such that for any pair $a,b\in X$ with $\operatorname{\mathsf{d}}(a,b)\geq R$, we have
#$\#\bigl{\{}g\in G\,|\,\operatorname{\mathsf{d}}(ga,a)\leq r\,\,\textup{and}\,\,d(gb,b)\leq r\bigr{\}}\leq N.$
Let $X$ be a hyperbolic space and $G$ be a group acting acylindrically on $X$ with a loxodromic element $g$. Some basic properties of this kind of actions can be found in [Osi16]. In this case, the elementary closure $EC(g)$ is the unique maximal virtually cyclic subgroup of $G$ that contains $g$ [DGO17, Lemma 6.5]. Moreover, $EC(g)$ has a subgroup of index at most $2$ that is a centralizer of some power of $g$ in $G$ [DGO17, Corollary 6.6].
In this paper, we will consider group actions with a large kernel, in which case the action cannot be acylindrical. As in [BBF21], an action $G\curvearrowright X$ is said to have acylindrical image if the image of $G$ in the isometry group of $X$ is acylindrical.
2.2. Projection axioms
In this section, we review the construction of a quasi-tree of spaces in [BBF15] with improvements from [BBFS19].
Let $\mathbf{Y}$ be a collection of geodesic metric spaces, and $\pi_{Y}(X)\subset Y$ be specified subsets whenever $X\neq Y$ are elements of $\mathbf{Y}$. Write $\operatorname{\mathsf{d}}^{\pi}_{Y}(X,Z)$ to mean $\operatorname{\mathsf{diam}}(\pi_{Y}(X)\cup\pi_{Y}(Z))$ for $X\neq Y\neq Z$. We say that $(\mathbf{Y},\{\pi_{Y}\})$ is a projection system with projection constant $\xi\geq 0$ if it satisfies the following projection axioms.
  1. (P0)
    $\operatorname{\mathsf{diam}}(\pi_{Y}(X))\leq\xi$ when $X\neq Y$; (Bounded projection)
  2. (P1)
    if $X,Y,Z$ are distinct and $\operatorname{\mathsf{d}}^{\pi}_{Y}(X,Z)>\xi$, then $\operatorname{\mathsf{d}}^{\pi}_{X}(Y,Z)\leq\xi$; (Behrstock inequality)
  3. (P2)
    for $X\neq Z$ the set
    $\{Y\in\mathbf{Y}\mid\operatorname{\mathsf{d}}^{\pi}_{Y}(X,Z)>\xi\}$
    is finite. (Finiteness)
Moreover, we say that $(\mathbf{Y},\{\pi_{Y}\})$ is a $G$projection system if a group $G$ acts on the set $\mathbf{Y}$ in such a way that every $g\in G$ acts as an isometry from $Y$ to $gY$ and the projections $\pi_{Y}$ are $G$–equivariant, that is, $\pi_{gY}(gX)=g\pi_{Y}(X)$.
If we replace (P1) with
  1. (P1)
    if $X,Y,Z$ are distinct and $\operatorname{\mathsf{d}}^{\pi}_{Y}(X,Z)>\xi$, then $\pi_{X}(Y)=\pi_{X}(Z)$,
then we say that $(\mathbf{Y},\{\pi_{Y}\})$ satisfies the strong projection axioms. While there are many natural situations where the projection axioms hold, the strong projection axioms are not as natural. However, we can modify the projections so that they do hold.
Theorem 2.1.
[BBFS19, Theorem 4.1] If $(\mathbf{Y},\{\pi_{Y}\})$ is a projection system with constant $\xi$, then there are projections $\{\pi^{\prime}_{Y}\}$ such that $(\mathbf{Y},\{\pi^{\prime}_{Y}\})$ satisfies the strong projections axioms with constant $\xi^{\prime}$, where $\pi^{\prime}_{Y}(X)$ and $\pi_{Y}(X)$ are apart from each other within a uniform Hausdorff distance $\epsilon$, and $\epsilon$ and $\xi^{\prime}$ only depend on $\xi$. Moreover, if $(\mathbf{Y},\{\pi_{Y}\})$ is a $G$–projection system, then $(\mathbf{Y},\{\pi^{\prime}_{Y}\})$ is still a $G$–projection system.
Let $\mathcal{C}_{K}\mathbf{Y}$ denote the space obtained from the disjoint union
$\bigsqcup_{Y\in\mathbf{Y}}Y$
by joining points in $\pi_{X}(Z)$ with points in $\pi_{Z}(X)$ by an edge of length one whenever $d_{Y}(X,Z)<K$ for all $Y\in\mathbf{Y}-\{X,Z\}$. When the spaces are graphs and projections are subgraphs, we can join just the vertices in these projections. Moreover, if $\mathbf{Y}$ is a $G$–projection system, then $G$ acts isometrically on $\mathcal{C}_{K}\mathbf{Y}$.
Theorem 2.2.
[BBF15, §4] If $(\mathbf{Y},\{\pi_{Y}\})$ satisfies the strong projection axioms with constant $\xi$, then for all $K>2\xi$
  • $\mathcal{C}_{K}\mathbf{Y}$ is hyperbolic if all $Y\in\mathbf{Y}$ are $\delta$-hyperbolic;
  • $\mathcal{C}_{K}\mathbf{Y}$ is a quasi-tree if all $Y\in\mathbf{Y}$ are quasi-trees with uniform QI constants.
There is a very useful distance formula in $\mathcal{C}_{K}\mathbf{Y}$. Let $X,Z\in\mathbf{Y}$ and $x\in X$, $z\in Z$. We define $\pi_{Y}(x)=\pi_{Y}(X)$ if $Y\neq X$ and define $\pi_{X}(x)=x$. Then define $\operatorname{\mathsf{d}}_{Y}(x,z)=\operatorname{\mathsf{diam}}(\pi_{Y}(x)\cup\pi_{Y}(z))$.
Theorem 2.3.
[BBFS19, Thm. 6.3] Let $(\mathbf{Y},\{\pi_{Y}\})$ satisfy the strong projection axioms with constant $\xi$. Let $x\in X$ and $z\in Z$ be two points of $\mathcal{C}(\mathbf{Y})$ with $X,Z\in\mathbf{Y}$. Then
$\frac{1}{4}\sum_{Y\in\mathbf{Y}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{Y}(x,z)\right\}\kern-3.01385pt\right\}_{K}\leq\operatorname{\mathsf{d}}_{\mathcal{C}_{K}\mathbf{Y}}(x,z)\leq 2\sum_{Y\in\mathbf{Y}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{Y}(x,z)\right\}\kern-3.01385pt\right\}_{K}+3K$
for all $K\geq 4\xi$.
Next we recall a theorem that allows us to pass projection axioms from a projection system to a collection of certain subspaces. Let $\mathbf{Y}$ be a collection of $\delta$–hyperbolic spaces and $(\mathbf{Y},\{\pi_{Y}\})$ be a projection system with constant $\xi$. For each $Y\in\mathbf{Y}$, let $\mathcal{A}_{Y}$ be a collection of quasi-geodesics in $Y$. Let $\mathcal{A}$ be the disjoint union of all $\mathcal{A}_{Y}$’s. We also make the following assumptions.
  • As a collection of quasi-geodesics, $\mathcal{A}$ has uniform coarse constants.
  • For $\alpha,\beta\in\mathcal{A}_{Y}$, we define $\pi_{\alpha}(\beta)$ to be the closest-point projection of $\beta$ to $\alpha$.
  • For $\alpha\in\mathcal{A}_{X}$ and $\beta\in\mathcal{A}_{Y}$ where $X\neq Y$, we define $\pi_{\alpha}(\beta)$ to be the closest point projection of $\pi_{X}(Y)$ to $\alpha$.
Theorem 2.4.
[BBF21, Thm. 4.17] For any $\theta>0$, there exists $\xi^{\prime}>0$, depending only on $\theta,\delta,\xi$ and coarse constants of $\mathcal{A}$, such that the following holds. If $\operatorname{\mathsf{diam}}(\pi_{\alpha}(\beta))\leq\theta$ whenever $\alpha$ and $\beta$ are distinct elements in the same $\mathcal{A}_{Y}$, then $(\mathcal{A},\{\pi_{\alpha}\})$ is a projection system with constant $\xi^{\prime}$.
2.3. (Relatively) hierarchically hyperbolic spaces
In this paper, we deal with (relatively) hierarchically hyperbolic spaces and (relatively) hierarchically hyperbolic groups. Coarsely speaking, a (relative) HHS is a pair $(\mathcal{X},\mathfrak{S})$, where $\mathcal{X}$ is a quasi-geodesic space and $\mathfrak{S}$ is an index set, with some extra structure. A full definition can be found in [BHS19, Definition 1.1,1.21]. Some important information from the definition is collected below.
[…]
  • For $U\sqsubsetneq V$ or $U\pitchfork V$, there is a uniformly bounded set $\rho^{U}_{V}\subset\mathcal{C}V$.
  • For $U\sqsubsetneq V$, there is a coarse map $\rho^{V}_{U}:\mathcal{C}V\to\mathcal{C}U$.
  • Whenever $V\sqsubseteq W$ and $W\perp U$, we require that $V\perp U$.
[…]
  • (Consistency) For every $x\in X$, the tuple $(\pi_{U}(x))_{U\in\mathfrak{S}}$ is $\kappa_{0}$–consistent (defined below). If $U\sqsubseteq V$, then $\operatorname{\mathsf{d}}_{W}(\rho^{U}_{W},\rho^{V}_{W})\leq\kappa_{0}$ whenever $W\in\mathfrak{S}$ satisfies either $V\sqsubsetneq W$ or $V\pitchfork W$ and $U\not\perp W$.
  • (Bounded geodesic image) There exists $E>0$ such that for all $W\in\mathfrak{S}$, all $V\in\mathfrak{S}_{W}^{\circ}$, and all $x,y\in\mathcal{X}$ so that some geodesic from $\pi_{W}(x)$ to $\pi_{W}(y)$ stays $E$–far from $\rho^{V}_{W}$, we have $\operatorname{\mathsf{d}}_{V}(\pi_{V}(x),\pi_{V}(y))\leq E$. We will refer this property as BGI in this paper.
Definition 2.5 ($\kappa$–consistent tuple).
For a number $\kappa\geq 0$, let $\vec{b}=(b_{U})_{U\in\mathfrak{S}}\in\prod_{U\in\mathfrak{S}}{2^{\mathcal{C}U}}$ be a tuple such that every set $b_{U}$ has diameter at most $\kappa$. We say that $\vec{b}$ is $\kappa$–consistent if
$\min\big{\{}\operatorname{\mathsf{d}}_{U}(b_{U},\rho^{V}_{U}),\operatorname{\mathsf{d}}_{V}(b_{V},\rho^{U}_{V})\big{\}}\leq\kappa\hskip 5.69054pt\text{ whenever }U\pitchfork V,\text{ and}$
[…]
In application, we always take $E$ to be the greatest constant in all coarseness from the above list (see [BHS19, Remark 1.6] for discussions on these constants). For the rest of this subsection, let $(\mathcal{X},\mathfrak{S})$ be a relative HHS.
Notation.
Given $x,y\in\mathcal{X}$, we write $\operatorname{\mathsf{d}}_{U}(x,y)$ to mean $\operatorname{\mathsf{d}}_{U}(\pi_{U}(x),\pi_{U}(y))$. If $U\pitchfork V$ or $U\sqsubsetneq V$, then we write $\operatorname{\mathsf{d}}_{V}(x,\rho^{U}_{V})$ to mean $\operatorname{\mathsf{d}}_{V}(\pi_{V}(x),\rho^{U}_{V})$.
Notation.
Given $A,B\geq 0$, we define a threshold function by
$\left\{\kern-3.01385pt\left\{A\right\}\kern-3.01385pt\right\}_{B}=\left\{\begin{array}[]{cl}A&\mbox{ if }A\geq B\\ 0&\mbox{ otherwise.}\end{array}\right.$
Given two functions $f,g:X\to\mathbb{R}$ and $A,B>0$, we write $f\preceq_{(A,B)}g$ to mean $f(x)\leq Ag(x)+B$ for any $x\in X$. We write $f\asymp_{(A,B)}g$ to mean $\frac{1}{A}f(x)-B\leq g(x)\leq Af(x)+B$ for any $x\in X$. Sometimes we omit the constants to mean that the inequality holds for some constants.
The powerful Masur–Minsky distance formula [MM00] shows that the distance between points in a mapping class group is coarsely the sum of the distances between the projections of these points to the curve graphs of all subsurfaces. Like mapping class groups, relative HHSes also satisfy a Masur–Minsky-style distance formula.
Theorem 2.6 (Distance formula).
[BHS19, Thm. 6.10] There exists $s_{0}$ such that for all $s\geq s_{0}$ there exists a constant $C>0$ such that for all $x,y\in\mathcal{X}$,
$\operatorname{\mathsf{d}}_{\mathcal{X}}(x,y)\asymp_{(C,C)}\sum_{W\in\mathfrak{S}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{s}.$
Closely related to the distance formula is the existence of hierarchy paths.
Definition 2.7 (Hierarchy path).
A $(D,D)$–quasi-geodesic $\gamma\subset\mathcal{X}$ is a $D$hierarchy path if $\pi_{U}(\gamma)$ is an unparametrized $(D,D)$–quasi-geodesic for each $U\in\mathfrak{S}$.
Theorem 2.8 (Existence of hierarchy paths).
[BHS19, Thm. 6.11] There exists $D_{0}$ such that any two points in $\mathcal{X}$ are joined by a $D_{0}$–hierarchy path.
Remark 2.9.
Let $\gamma$ be a $D$–hierarchy path connecting $x$ and $y$. By the construction of hierarchy paths in [BHS19], $\pi_{U}(\gamma)$ is contained in the $D$–neighborhood of a geodesic connecting $\pi_{U}(x)$ and $\pi_{U}(y)$. If $\mathcal{C}U$ is hyperbolic, this is easy to see from Morse Lemma. Otherwise, this deserves its own mention.
There is an important class of subspaces in relative HHSes. We will consider them in Section 4.
Definition 2.10 (Standard product region, standard nested region, standard orthogonal region).
Given $U\in\mathfrak{S}$. Fix $\kappa\geq\kappa_{0}$. Let $\mathbf{F}_{U}$ be the set of $\kappa$–consistent tuples in $\prod_{V\in\mathfrak{S}_{U}}2^{\mathcal{C}V}$. Let $\mathbf{E}_{U}$ be the set of $\kappa$–consistent tuples in $\prod_{V\in\mathfrak{S}_{U}^{\perp}-\{A\}}2^{\mathcal{C}V}$. Let $\mathbf{P}_{U}=\mathbf{F}_{U}\times\mathbf{E}_{U}$. We can define a coarse map $\phi_{U}:\mathbf{P}_{U}\to\mathcal{X}$ as follows.
For each $(\vec{a},\vec{b})\in\mathbf{F}_{U}\times\mathbf{E}_{U}$, and each $V\in\mathfrak{S}$, define the coordinate $(\phi_{U}(\vec{a},\vec{b}))_{V}$ as follows. If $V\sqsubseteq U$, then $(\phi_{U}(\vec{a},\vec{b}))_{V}=a_{V}$. If $V\perp U$, then $(\phi_{U}(\vec{a},\vec{b}))_{V}=b_{V}$. If $V\pitchfork U$, then $(\phi_{U}(\vec{a},\vec{b}))_{V}=\rho^{U}_{V}$. Finally, if $U\sqsubsetneq V$, let $(\phi_{U}(\vec{a},\vec{b}))_{V}=\rho^{U}_{V}$. We can check that the tuple $\phi_{U}(\vec{a},\vec{b})$ is $\kappa$–consistent, and thus the realization theorem [BHS19, Theorem 3.1] supplies the map $\phi_{U}:\mathbf{P}_{U}\to\mathcal{X}$ (see [BHS19, §5B] for more details).
For convenience, we do not distinguish between $\mathbf{P}_{U}$ and its image in $\mathcal{X}$. We call $\mathbf{P}_{U}$ the standard product region. By choosing any copy of $\mathbf{F}_{U}$ in the direct product, $\phi_{U}$ restricts to a coarse map $\phi^{\sqsubseteq}:\mathbf{F}_{U}\to\mathcal{X}$. We also define $\phi^{\perp}:\mathbf{E}_{U}\to\mathcal{X}$ in the same way. We call $\mathbf{F}_{U}$ and $\mathbf{E}_{U}$ the standard nested region and the standard orthogonal region, respectively.
Remark 2.11.
By definition, $\mathbf{F}_{U}$, $\mathbf{E}_{U}$ and $\mathbf{P}_{U}$ depend on the constant $\kappa$. In this paper, we simply fix any $\kappa\geq\kappa_{0}$ and do not mention it again.
It is known that $(\mathbf{F}_{U},\mathfrak{S}_{U})$, $(\mathbf{E}_{U},\mathfrak{S}_{U}^{\perp})$ are both relatively hierarchically hyperbolic. By definition of $\mathbf{F}_{U}$, $\mathbf{E}_{U}$ and $\mathbf{P}_{U}$, there are natural retractions from $\mathcal{X}$ to these subspaces. We call such a map a gate map. Take $\mathbf{F}_{U}$ for example. We denote the gate map to $\mathbf{F}_{U}$ by $\mathfrak{g}_{\mathbf{F}_{U}}$. Note that $\pi_{V}(\mathbf{F}_{U})\subset\mathcal{C}V$ is quasi-convex since $\pi_{V}(\mathbf{F}_{U})$ is coarsely $\pi_{V}(\mathcal{X})$. Furthermore, for all $x\in\mathcal{X}$ and all $V\in\mathfrak{S}$ such that $\mathcal{C}V$ is hyperbolic, $\pi_{V}(\mathfrak{g}_{\mathbf{F}_{U}}(x))$ uniformly coarsely coincides with the closest point projection of $\pi_{V}(x)$ to $\pi_{V}(\mathbf{F}_{U})$. In fact, gate maps can be defined for all “hierarchically quasi-convex” subspaces, which is a larger class of subspaces of relative HHSes (see [BHS19, §5] for HHSes).
For any (relative) HHS $(\mathcal{X},\mathfrak{S})$, an automorphism is roughly speaking a bijection from $(\mathcal{X},\mathfrak{S})$ to itself that preserves its (relative) HHS structure [BHS19, §1G]. All automorphisms of $(\mathcal{X},\mathfrak{S})$ form a group $\operatorname{Aut}(\mathfrak{S})$, which we call the automorphism group of $(\mathcal{X},\mathfrak{S})$.
Definition 2.12 ((Relatively) hierarchically hyperbolic groups).
A finitely generated group $G$ is (relatively) hierarchically hyperbolic if there exists a (relatively) hierarchically hyperbolic space $(\mathcal{X},\mathfrak{S})$ and an action $G\to\operatorname{Aut}(\mathfrak{S})$ such that the action $G\curvearrowright X$ is metrically proper and cobounded, and such that the induced action on $\mathfrak{S}$ is cofinite.
Note that if $G$ is (relatively) hierarchically hyperbolic by virtue of its action on the (relatively) hierarchically hyperbolic space $(\mathcal{X},\mathfrak{S})$, then $(G,\mathfrak{S})$ is a (relatively) hierarchically hyperbolic structure with respect to any word-metric on $G$.
Let $\operatorname{Aut}(\mathfrak{S};V)$ be the group of automorphisms $g\in\operatorname{Aut}(\mathfrak{S})$ such that $g\cdot V=V$. Then there is a restriction homomorphism $\theta_{V}:\operatorname{Aut}(\mathfrak{S};V)\to\operatorname{Aut}(\mathfrak{S}_{V})$ defined as follows. Given $g\in\operatorname{Aut}(\mathfrak{S};V)$, let $\theta_{V}(g)$ act like $g$ on the substructure $\mathfrak{S}_{V}$. For a group $G<\text{Aut}(\mathfrak{S})$, we write $\operatorname{Stab}_{G}(V)$ to mean $G\cap\operatorname{Aut}(\mathfrak{S};V)$ and write $G_{V}$ to mean the image of $\operatorname{Stab}_{G}(V)$ under $\theta_{V}$.
For many HHGs (for example, the case of mapping class groups), every $G_{V}$ acts acylindrically on $\mathcal{C}V$. However, not all HHGs have this property [DHS20].
Definition 2.13 (colorability).
Let $(G,\mathfrak{S})$ be a relative HHG. Let $\mathfrak{S}^{\prime}\subset\mathfrak{S}$ be a $G$–invariant subset. We say $\mathfrak{S}^{\prime}$ is colorable if, $\mathfrak{S}^{\prime}$ admits a decomposition $\mathfrak{S}^{\prime}=\bigsqcup_{i=1}^{\chi}\mathfrak{S}^{\prime}_{i}$ into finitely many $G$–invariant families $\mathfrak{S}^{\prime}_{i}$ such that any two domains in the same family are transverse. Such a decomposition is called a coloring of $\mathfrak{S}^{\prime}$. We say a relative HHG $(G,\mathfrak{S})$ is colorable if $\mathfrak{S}$ is colorable.
The notion of colorability is formalized in [DMS23, HP22]. There are many classes of (virtually) colorable HHGs, as listed in the above papers. In particular, a coloring is constructed for (a finite-index subgroup of) a mapping class group in [BBF15, §5], from which the notion comes. However, one cannot expect that all HHGs are virtually colorable [Hag23]. Nevertheless, [HP22, Prop. 3.2] provides a sufficient condition for an HHG to be virtually colorable.
Remark 2.14.
In this paper, we only concern unbounded domains, i.e. domains with unbounded associated hyperbolic spaces. As an abuse of terminology, we say a relative HHG $(G,\mathfrak{S})$ is colorable if the collection of unbounded domains is colorable.
3. Thick distance formula
In this section, we will prove a thick distance formula that is similar to [BBF21, Theorem 4.13]. This allows us to estimate the distance in a relative HHS by counting only “thick” segments of a hierarchy path instead of the whole hierarchy path. The readers should be aware that the definitions in this section are different from those [BBF21]. In particular, we do not have tight geodesics in a general HHG.
[…]
Lemma 3.1.
Given $x,y\in\mathcal{X}$ and $U\in\operatorname{Rel}_{T}(x,y)$, there exist at most two domains $V_{1},V_{2}\in\operatorname{Rel}_{T}(x,y)$ such that $U\in\operatorname{Rel}_{T}^{m}(V_{i};x,y)$ for $i=1,2$.
Proof.
[…]
By [BHS19, Proposition 2.8], any set of pairwise transverse elements in $\operatorname{Rel}_{T}(x,y)$ has a total order $<$, obtained by setting $U<V$ whenever $\operatorname{\mathsf{d}}_{U}(y,\rho^{V}_{U})\leq E$. We assume that $V_{1}<V_{2}<V_{3}$.
On the one hand, $\operatorname{\mathsf{d}}_{V_{2}}(\rho^{V_{1}}_{V_{2}},\rho^{V_{3}}_{V_{2}})\geq\operatorname{\mathsf{d}}_{V_{2}}(x,y)-2E>T-2E$ by the triangle inequality. On the other hand, $\operatorname{\mathsf{d}}_{V_{2}}(\rho^{V_{1}}_{V_{2}},\rho^{U}_{V_{2}})\leq\kappa_{0}$ and $\operatorname{\mathsf{d}}_{V_{2}}(\rho^{V_{3}}_{V_{2}},\rho^{U}_{V_{2}})\leq\kappa_{0}$ by consistency, which gives $\operatorname{\mathsf{d}}_{V_{2}}(\rho^{V_{1}}_{V_{2}},\rho^{V_{3}}_{V_{2}})\leq 2\kappa_{0}<T-2E$. This gives a contradiction. ∎
Definition 3.2 ($T$–thickness).
Given $\mathfrak{S}^{\prime}\subset\mathfrak{S}$, we say a pair of points $(x,y)\in\mathcal{X}\times\mathcal{X}$ is $T$–thick for $\mathfrak{S}^{\prime}$ if $\operatorname{\mathsf{diam}}(\pi_{U}(x)\cup\pi_{U}(y))\leq T$ for all $U\in\mathfrak{S}^{\prime}$. We define $\mathcal{P}_{T}(\mathfrak{S}^{\prime})$ to be the set of all $T$–thick pairs of points for $\mathfrak{S}^{\prime}$. If $\mathfrak{S}^{\prime}=\{U\}$, we also say $(x,y)$ is $T$–thick for $U$ and write $(x,y)\in\mathcal{P}_{T}(U)$
Note that $(x,y)\in\mathcal{P}_{T}(U)$ if and only if $U\notin\operatorname{Rel}_{T}(x,y)$. Also note that $(x,y)\in\mathcal{P}_{T}(\mathfrak{S}_{V}^{\circ})$ if and only if $\operatorname{Rel}_{T}(V;x,y)=\emptyset$.
Lemma 3.3.
Let $D_{0}$ be the constant provided by Theorem 2.8. For any $x,y\in\mathcal{X}$, let $\gamma$ be a $D_{0}$–hierarchy path between $x,y$. Given any $U\in\mathfrak{S}$ and any $x^{\prime},y^{\prime}\in\gamma$, then
$\operatorname{\mathsf{d}}_{U}(x^{\prime},y^{\prime})\leq\operatorname{\mathsf{d}}_{U}(x,y)+2D_{0}.$
In particular, if $(x,y)\in\mathcal{P}_{T}(U)$, then $(x^{\prime},y^{\prime})\in\mathcal{P}_{T+2D_{0}}(U)$.
Proof.
By Remark 2.9, $\pi_{U}(\gamma)$ lies in the $D_{0}$–neighborhood of a geodesic connecting $\pi_{U}(x)$ and $\pi_{U}(y)$. The conclusion then follows from the triangle inequality. ∎
Notation.
Let $\hat{T}=T+2D_{0}$ and $\check{T}=T-2D_{0}$.
Notation.
Given two points $x,y$ in a hyperbolic space, we write $[x,y]$ to mean a geodesic segment between $x,y$, which is coarsely unique. For an interval $I$ or a path $\gamma$, we write $I^{-},I^{+}$ or $\gamma^{-},\gamma^{+}$ to mean their endpoints.
Definition 3.4 ($(T,R)$–thick distance).
Fix sufficiently large constants $T,R$. Let $\gamma$ be a $D_{0}$–hierarchy path between $x$ and $y$. Let $\gamma_{1},\dots,\gamma_{n}\subset\gamma$ be disjoint subpaths occurring in this order such that $(\gamma_{i}^{-},\gamma_{i}^{+})\in\mathcal{P}_{T}(\mathfrak{S}_{V}^{\circ})$ for each $i$.
The $(T,R)$–thick distance in $V$ is denoted by $\operatorname{\mathsf{d}}_{V}^{T,R}(x,y)$ and is defined to be the supremum of $\sum_{i=1}^{n}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{V}(\gamma_{i}^{-},\gamma_{i}^{+})\right\}\kern-3.01385pt\right\}_{R}$ over all such choices for $\gamma_{i}$’s, and for all $D_{0}$–hierarchy paths from $x$ to $y$.
[…]
Lemma 3.5.
Fix constants $T,R>100E$. For any $x,y\in\mathcal{X}$ and $W\in\mathfrak{S}$, we have
$\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}\leq\operatorname{\mathsf{d}}^{T,R}_{W}(x,y)+(6E+2R)|\operatorname{Rel}_{\check{T}}^{m}(W;x,y)|.$
Proof.
If $\operatorname{Rel}_{\check{T}}^{m}(W)=\emptyset$, then $(x,y)\in\mathcal{P}_{T}(\mathfrak{S}_{W}^{\circ})$. Thus, both sides of the above inequality are equal.
[…]
$s_{V}^{-}:=\inf\{s\in I\mid\exists U\sqsubseteq V\text{such that}\operatorname{\mathsf{d}}_{W}(\gamma(s),\rho^{U}_{W})\leq 2E\},$
$s_{V}^{+}:=\sup\{s\in I\mid\exists U\sqsubseteq V\text{such that}\operatorname{\mathsf{d}}_{W}(\gamma(s),\rho^{U}_{W})\leq 2E\}.$
For any $U\sqsubseteq V\sqsubsetneq W$, we know that $\operatorname{\mathsf{d}}_{W}(\rho^{U}_{W},\rho^{V}_{W})\leq\kappa_{0}$ by consistency. Thus,
$\operatorname{\mathsf{d}}_{W}(\gamma(s_{V}^{-}),\rho^{V}_{W}))\leq 2E+\kappa_{0}\leq 3E,$
$\operatorname{\mathsf{d}}_{W}(\gamma(s_{V}^{+}),\rho^{V}_{W}))\leq 2E+\kappa_{0}\leq 3E.$
Therefore,
$\operatorname{\mathsf{d}}_{W}(\gamma(s_{V}^{-}),\gamma(s_{V}^{+}))\leq\operatorname{\mathsf{d}}_{W}(\gamma(s_{V}^{-}),\rho^{V}_{W}))+\operatorname{\mathsf{d}}_{W}(\gamma(s_{V}^{+}),\rho^{V}_{W}))\leq 6E.$
Let $J_{0},\dots,J_{n}$ be the collection of maximal intervals in $I-\bigcup_{V\in\operatorname{Rel}_{\check{T}}^{m}(W)}(s_{V}^{-},s_{V}^{+})$. Note that $n\leq|\operatorname{Rel}_{\check{T}}^{m}(W)|$. Now we are going to prove that $(\gamma(J_{i}^{-}),\gamma(J_{i}^{+}))\in\mathcal{P}_{T}(\mathfrak{S}_{W}^{\circ})$.
On the one hand, $\operatorname{\mathsf{d}}_{W}(\rho_{W}^{U},\gamma(J_{i}))\geq 2E$ for any $U\in\operatorname{Rel}_{\check{T}}(W)$ by definition of $J_{i}$. Then by Morse Lemma,
$\operatorname{\mathsf{d}}_{W}(\rho_{W}^{U},[\gamma(J_{i}^{-}),\gamma(J_{i}^{+})])\geq\operatorname{\mathsf{d}}_{W}(\rho_{W}^{U},\gamma(J_{i}))-E\geq E.$
Therefore, $\operatorname{\mathsf{d}}_{U}(\gamma(J_{i}^{-}),\gamma(J_{i}^{+}))\leq E<T$ by BGI. On the other hand, $(x,y)$ is $\check{T}$–thick for $\mathfrak{S}_{W}^{\circ}-\operatorname{Rel}_{\check{T}}(W)$ by definition. It follows from Lemma 3.3 that $(\gamma(J_{i}^{-}),\gamma(J_{i}^{+}))$ is $T$–thick for $\mathfrak{S}_{W}^{\circ}-\operatorname{Rel}_{\check{T}}(W)$. In summary, $(\gamma(J_{i}^{-}),\gamma(J_{i}^{+}))\in\mathcal{P}_{T}(\mathfrak{S}_{W}^{\circ})$.
Finally, we estimate that
$\operatorname{\mathsf{d}}_{W}(x,y)$ $\leq\sum_{i=0}^{n}\operatorname{\mathsf{d}}_{W}(\gamma(J_{i}^{+}),\gamma(J_{i}^{-}))+6E|\operatorname{Rel}_{\check{T}}^{m}(W)|$
$\leq\operatorname{\mathsf{d}}^{T,R}_{W}(x,y)+R(n+1)+6E|\operatorname{Rel}_{\check{T}}^{m}(W)|$
$\leq\operatorname{\mathsf{d}}^{T,R}_{W}(x,y)+(6E+2R)|\operatorname{Rel}_{\check{T}}^{m}(W)|.$
Let $S$ denote the unique maximal domain in $\mathfrak{S}$. Recall that the level $\ell(S)$ of $S$ is equal to the complexity of $(\mathcal{X},\mathfrak{S})$.
Theorem 3.6.
Fix the constants $T,R$ with $\check{T}\geq R>100E$. Let $x,y\in\mathcal{X}$. Then, for each $n$
$\sum_{\ell(W)\leq n}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}\leq\sum_{\ell(W)=n}\operatorname{\mathsf{d}}^{T,R}_{W}(x,y)+7\sum_{\ell(W)<n}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}.$
Note that each sum is over finitely many $W$ since there are only finitely many $W$ such that $\operatorname{\mathsf{d}}_{W}(x,y)\geq R$ for given $x,y$ by the distance formula (Theorem 2.6).
Proof.
If $\ell(W)=n$ then by Lemma 3.5,
$\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}$ $\leq\operatorname{\mathsf{d}}^{T,R}_{W}(x,y)+(6E+2R)|\operatorname{Rel}^{m}_{\check{T}}(W)|$
$\leq\operatorname{\mathsf{d}}^{T,R}_{W}(x,y)+(6E+2R)\sum_{V\in\operatorname{Rel}^{m}_{\check{T}}(W)}\frac{\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{V}(x,y)\right\}\kern-3.01385pt\right\}_{\check{T}}}{\check{T}}$
$\leq\operatorname{\mathsf{d}}^{T,R}_{W}(x,y)+3\sum_{V\in\operatorname{Rel}^{m}_{\check{T}}(W)}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{V}(x,y)\right\}\kern-3.01385pt\right\}_{\check{T}}$
$\leq\operatorname{\mathsf{d}}^{T,R}_{W}(x,y)+3\sum_{V\in\operatorname{Rel}^{m}_{\check{T}}(W)}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{V}(x,y)\right\}\kern-3.01385pt\right\}_{R}.$
By Lemma 3.1, any $V$ appears in at most two $\operatorname{Rel}^{m}_{\check{T}}(W)$. Therefore, if we sum up the left hand side over all $W$ with $\ell(W)=n$, we have
$\sum_{\ell(W)=n}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}\leq\sum_{\ell(W)=n}\operatorname{\mathsf{d}}^{T,R}_{W}(x,y)+6\sum_{\ell(W)<n}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}.$
Adding $\sum_{\ell(W)<n}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}$ to both sides gives the desirable inequality. ∎
Corollary 3.7.
Fix the constants $T,R$ with $\check{T}\geq R>100E$. Let $x,y\in\mathcal{X}$. Then
$\frac{1}{D_{0}}\sum_{W\in\mathfrak{S}}\operatorname{\mathsf{d}}_{W}^{T,R}(x,y)-D_{0}\leq\sum_{W\in\mathfrak{S}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}\leq 7^{\ell(S)-1}\sum_{W\in\mathfrak{S}}\operatorname{\mathsf{d}}_{W}^{T,R}(x,y).$
Proof.
The first inequality is trivial since $\frac{1}{D_{0}}\operatorname{\mathsf{d}}_{W}^{T,R}(x,y)-D_{0}\leq\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}$ for all $W$. By inductively applying Theorem 3.6, with base case $n=\ell(S)$, we have
$\sum_{W\in\mathfrak{S}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}\leq 7^{\ell(S)-n}\left(\sum_{n\leq\ell(W)\leq\ell(S)}\operatorname{\mathsf{d}}^{T,R}_{W}(x,y)+7\sum_{\ell(W)<n}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{W}(x,y)\right\}\kern-3.01385pt\right\}_{R}\right).$
When $n=1$, the last term on the right is zero, and the result follows. ∎
Combining the distance formula (Theorem 2.6) with Corollary 3.7, we obtain our thick distance formula.
Theorem 3.8 (Thick distance formula).
There exists $R_{0}$ such that for all $T,R$ with $\check{T}\geq R>R_{0}$, there exists a constant $L>0$ such that for all $x,y\in\mathcal{X}$,
$\operatorname{\mathsf{d}}_{\mathcal{X}}(x,y)\asymp_{(L,L)}\sum_{W\in\mathfrak{S}}\operatorname{\mathsf{d}}_{W}^{T,R}(x,y).$
4. Estimation of thick distance via quasi-axes
Our proof in this section is inspired by [NY23]. The main technique in the proof of [NY23, Lemma 5.5] that is different from [BBF21] is the use of the Extension Lemma (see Lemma 4.1 below). This lemma is widely applicable so it benefits us to deal with HHGs. In the original statement of [NY23, Lem. 4.13], it is required that the group action is cobounded, but it does not matter if we relax this condition. The readers could compare it with [BBF21, Theorem 4.19] for mapping class groups.
Lemma 4.1 (Extension Lemma).
[NY23, Lem. 4.13] Let $H$ be a non-elementary group acting acylindrically on a $\delta$–hyperbolic space $Y$. Fix a base point $o\in Y$. There exists a set $F\subset H$ of three loxodromic elements and constants $\lambda\geq 1,c\geq 0$ with the following property.
For any $h\in H$ there exists $f\in F$ such that $hf$ is a loxodromic element and the bi-infinite path
$\gamma_{h}=\bigcup_{i\in\mathbb{Z}}(hf)^{i}([o,ho][ho,hfo])$
is a $(\lambda,c)$–quasi-geodesic.
Let $(\mathcal{X},\mathfrak{S})$ be a relative HHS with a coarse constant $E$, and let $G$ be a relative HHG by virtue of its action on $\mathcal{X}$. Let $V\in\mathfrak{S}$ such that $V$ has hyperbolicity, cobounded nested region and acylindrical image.
Lemma 4.2 (Extension of thick segments).
There exist constants $\lambda\geq 1,c\geq 0,B\geq 0$ such that the following holds. For any $T,R>0$, there exists a $G_{V}$–finite collection $\mathcal{A}_{V}=\mathcal{A}_{V}^{T,R}$ of $(\lambda,c)$–quasi-axes in $\mathcal{C}V$ such that for any pair of points $(x,y)\in\mathcal{P}_{\max\{\hat{T},R\}}(\mathfrak{S}_{V})$, there exists $\gamma\in\mathcal{A}_{V}$ such that $[\pi_{V}(x),\pi_{V}(y)]\subset\mathcal{N}_{B}(\gamma)$.
Proof.
Fix a base point $o\in\mathbf{F}_{V}$ and project it to a base point in $\mathcal{C}V$. For the action $G_{V}\curvearrowright\mathcal{C}V$, Lemma 4.1 provides a finite set $F\subset G_{V}$ and constants $\lambda\geq 1,c\geq 0$. Recall that $(\mathbf{F}_{V},\mathfrak{S}_{V})$ is a relative HHS. $\mathbf{F}_{V}$ is proper because $\mathcal{X}$ is proper.
Since $G_{V}$ acts coboundedly on $\mathbf{F}_{V}$, there exists $\epsilon>0$ such that $\mathbf{F}_{V}$ is covered by the $G_{V}$–translates of any $\epsilon$–ball. Let $T^{\prime}=\max\{\hat{T},R\}$. By Theorem 2.6, there exists $r>0$, depending only on $E,T^{\prime},\epsilon$, such that the distance between any pair of points in $\mathbf{F}_{V}$ that is $(T^{\prime}+\epsilon)$–thick for $\mathfrak{S}_{V}$ is bounded above by $r$. Fix any base point $o\in\mathbf{F}_{V}$. Since $\mathbf{F}_{V}$ is proper, there exists a finite subset $S\subset G_{V}$ such that $\mathcal{N}_{r+\epsilon}(o)$ is covered by $\bigcup_{s\in S}s\cdot\mathcal{N}_{\epsilon}(o)$.
Lemma 4.1 tells us that for each $s\in S$, there exists $f\in F$ such that $sf$ is a loxodromic element acting on $\mathcal{C}V$. Let $\mathcal{A}_{V}=\mathcal{A}_{V}^{T,R}$ be the collection of $G_{V}$–translates of the $(\lambda,c)$–quasi-axes provided for all $s\in S$ by Lemma 4.1.
Now we verify that $\mathcal{A}_{V}$ meets our requirements. Let $(x,y)\in\mathcal{P}_{T^{\prime}}(\mathfrak{S}_{V})$. We can choose $h\in G_{V}$ such that $\operatorname{\mathsf{d}}_{\mathbf{F}_{V}}(x,ho)<\epsilon$. Then $\operatorname{\mathsf{d}}_{\mathbf{F}_{V}}(o,h^{-1}y)\leq\operatorname{\mathsf{d}}_{\mathbf{F}_{V}}(x,y)+\operatorname{\mathsf{d}}_{\mathbf{F}_{V}}(x,ho)<r+\epsilon$. By our choice of $S$, there exists $s\in S$ such that $\operatorname{\mathsf{d}}_{\mathbf{F}_{V}}(h^{-1}y,so)<\epsilon$. Thus, we have
$\operatorname{\mathsf{d}}_{V}(x,ho)$ $<E\epsilon+E,$
$\operatorname{\mathsf{d}}_{V}(y,hso)$ $<E\epsilon+E,$
because $\pi_{V}$ is $E$–coarsely Lipschitz. Since $\mathcal{C}V$ is $E$–hyperbolic, we can find $B>0$ by fellow-traveller property such that $[\pi_{V}(x),\pi_{V}(y)]\subset\mathcal{N}_{B}([h\cdot\pi_{V}(o),hs\cdot\pi_{V}(o)])$. By construction, $[h\cdot\pi_{V}(o),hs\cdot\pi_{V}(o)]$ is contained in some $\gamma\in\mathcal{A}_{V}$ so we are done. ∎
Notation.
Assume that $\gamma$ is a quasi-geodesic in a hyperbolic space $Y$. We write $\pi_{\gamma}:Y\to\gamma$ to mean the closest point projection. For $x,y\in Y$, we write $\operatorname{\mathsf{d}}_{\gamma}(x,y)$ to mean $\operatorname{\mathsf{diam}}(\pi_{\gamma}(x)\cup\pi_{\gamma}(y))$.
Notation.
Assume that $\gamma$ is a quasi-geodesic in $\mathcal{C}V$. We write $\pi^{\mathcal{X}}_{\gamma}$ to mean $\pi_{\gamma}\circ\pi_{V}$. For $x,y\in\mathcal{X}$, we write $\operatorname{\mathsf{d}}^{\mathcal{X}}_{\gamma}(x,y)$ to mean $\operatorname{\mathsf{diam}}(\pi^{\mathcal{X}}_{\gamma}(x)\cup\pi^{\mathcal{X}}_{\gamma}(y))$.
The following lemma is a well-known fact about $\delta$–hyperbolic spaces so we omit the proof.
Lemma 4.3.
Let $\gamma$ and $\alpha$ be two $(\lambda,c)$–quasi-geodesics in a $\delta$–hyperbolic space. Then for any $B>0$, there exists a constant $C=C(\lambda,c,B,\delta)>0$ such that
$\operatorname{\mathsf{d}}_{\gamma}(\alpha^{-},\alpha^{+})\geq\operatorname{\mathsf{diam}}(\alpha\cap\mathcal{N}_{B}(\gamma))-C.$
The main result of this section is the following estimation that generalizes [BBF21, Propostion 4.18].
Proposition 4.4.
For any $K>0$, there exists $R>0$ such that the following holds. Given any $T>0$, let $\mathcal{A}_{V}=\mathcal{A}_{V}^{T,R}$ be the collection of $(\lambda,c)$–quasi-axes provided by Lemma 4.2. Then for any two points $x,y\in\mathcal{X}$,
$\operatorname{\mathsf{d}}_{V}^{T,R}(x,y)\leq 2(D_{0}+1)\sum_{\gamma\in\mathcal{A}_{V}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}^{\mathcal{X}}_{\gamma}(x,y)\right\}\kern-3.01385pt\right\}_{K}$
where $D_{0}$ is the constant provided by Theorem 2.8.
Proof.
Let $C=C(\lambda+D_{0},c+D_{0},B,E)$ be the constant provided by Lemma 4.3. Let $R>2D_{0}(C+1)+K$. We will show that projections to quasi-axes $\mathcal{A}_{V}=\mathcal{A}_{V}^{T,R}$ bound the $(T,R)$–thick distance in $V$ from above.
For any two points $x,y\in\mathcal{X}$, let $\beta$ be a $(D_{0},D_{0})$–hierarchy path connecting $x$ and $y$ realizing $\operatorname{\mathsf{d}}_{V}^{T,R}(x,y)$. Let $\{\alpha_{1},\dots,\alpha_{n}\}$ be the collection of disjoint subpaths of $\beta$ with $\operatorname{\mathsf{d}}_{V}(\alpha_{i}^{-},\alpha_{i}^{+})\geq R$ and $(\alpha_{i}^{-},\alpha_{i}^{+})\in\mathcal{P}_{T}(\mathfrak{S}_{V}^{\circ})$ such that
$\operatorname{\mathsf{d}}^{T,R}_{V}(x,y)=\sum_{i=1}^{n}\operatorname{\mathsf{d}}_{V}(\alpha_{i}^{-},\alpha_{i}^{+}).$
By definition of gate maps, $\pi_{V}(z)$ is coarsely $\pi_{V}(\mathfrak{g}_{\mathbf{F}_{V}}(z))$ for any $z\in\mathcal{X}$. By substituting with the image under the gate map $\mathfrak{g}_{\mathbf{F}_{V}}$, we can assume that $\alpha_{i}$ is contained in $\mathbf{F}_{V}$ up to a bounded error of the thick distance in $V$. We divide each $\alpha_{i}$ into several consecutive subpaths $\{\tilde{\alpha}_{i,j}~{}|~{}1\leq j\leq m_{i}\}$ with $\operatorname{\mathsf{d}}_{V}(\tilde{\alpha}_{i,j}^{-},\tilde{\alpha}_{i,j}^{+})=R$ for $j=1,\dots,m_{i}-1$ and $\operatorname{\mathsf{d}}_{V}(\tilde{\alpha}_{i,m_{i}}^{-},\tilde{\alpha}_{i,m_{i}}^{+})\leq R$. By Lemma 3.3, we already know that $(\tilde{\alpha}_{i,j}^{-},\tilde{\alpha}_{i,j}^{+})\in\mathcal{P}_{\hat{T}}(\mathfrak{S}_{V}^{\circ})$ for every pair $(i,j)$. Thus,
$\operatorname{\mathsf{d}}_{U}(\tilde{\alpha}_{i,j}^{-},\tilde{\alpha}_{i,j}^{+})\leq\max\{\hat{T},R\}$
for all $U\in\mathfrak{S}_{V}$. By Lemma 4.2, there exists $\gamma_{i,j}\in\mathcal{A}_{V}$ such that $\pi_{U}(\tilde{\alpha}_{i,j})\subset\mathcal{N}_{B}(\gamma_{i,j})$ (with an increased $B$ by a uniform constant), which yields
$\operatorname{\mathsf{diam}}(\pi_{V}(\alpha_{i})\cap\mathcal{N}_{B}(\gamma_{i,j}))\geq R.$
Let $\mathcal{A}_{V}^{\prime}$ be the collection of all distinct $\gamma_{i,j}$’s. We see that
$\pi_{V}(\alpha_{i}\backslash\tilde{\alpha}_{i,m_{i}})\subset\bigcup_{\gamma\in\mathcal{A}_{V}^{\prime}}\pi_{V}(\alpha_{i})\cap\mathcal{N}_{B}(\gamma).$
Thus, we have
$\operatorname{\mathsf{d}}_{V}(\alpha_{i}^{-},\alpha_{i}^{+})$ $\leq\sum_{\gamma\in\mathcal{A}_{V}^{\prime}}\operatorname{\mathsf{diam}}(\pi_{V}(\alpha_{i})\cap\mathcal{N}_{B}(\gamma))+R$
$\leq 2\sum_{\gamma\in\mathcal{A}_{V}^{\prime}}\operatorname{\mathsf{diam}}(\pi_{V}(\alpha_{i})\cap\mathcal{N}_{B}(\gamma)).$
Summing up from $i=1$ to $n$ yields that
(1) $\operatorname{\mathsf{d}}^{T,R}_{V}(x,y)\leq 2\sum_{\gamma\in\mathcal{A}_{V}^{\prime}}(D_{0}\operatorname{\mathsf{diam}}(\pi_{V}(\beta)\cap\mathcal{N}_{B}(\gamma))+D_{0}).$
Note that $R>2D_{0}(C+1)+K$. Thus, Lemma 4.3 tells us that $\operatorname{\mathsf{d}}^{\mathcal{X}}_{\gamma}(x,y)\geq\operatorname{\mathsf{diam}}(\pi_{V}(\beta)\cap\mathcal{N}_{B}(\gamma))-C\geq R-C>D_{0}(C+1)+K$ for each $\gamma\in\mathcal{A}_{V}^{\prime}$. We now estimate by Lemma 4.3 that
$D_{0}\operatorname{\mathsf{diam}}(\pi_{V}(\beta)\cap\mathcal{N}_{B}(\gamma))+D_{0}$ $\leq D_{0}(\operatorname{\mathsf{d}}^{\mathcal{X}}_{\gamma}(x,y)+C)+D_{0}$
$<(D_{0}+1)\operatorname{\mathsf{d}}^{\mathcal{X}}_{\gamma}(x,y).$
Combining it with Equation (1), we obtain that
$\operatorname{\mathsf{d}}^{T,R}_{V}(x,y)$ $\leq 2(D_{0}+1)\sum_{\gamma\in\mathcal{A}_{V}^{\prime}}\operatorname{\mathsf{d}}^{\mathcal{X}}_{\gamma}(x,y)$
$=2(D_{0}+1)\sum_{\gamma\in\mathcal{A}_{V}^{\prime}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}^{\mathcal{X}}_{\gamma}(x,y)\right\}\kern-3.01385pt\right\}_{K}$
$\leq 2(D_{0}+1)\sum_{\gamma\in\mathcal{A}_{V}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}^{\mathcal{X}}_{\gamma}(x,y)\right\}\kern-3.01385pt\right\}_{K}.$
5. Construction of quasi-trees
This section devotes to the proof of Theorem 1.3. Let $(G,\mathfrak{S})$ be a relative HHG that is virtually colorable and assume that every domain in $\mathfrak{S}$ is of Type I or Type II. The index set $\mathfrak{S}$ admits a $G$–invariant decomposition $\mathfrak{S}=\mathfrak{S}^{I}\bigsqcup\mathfrak{S}^{II}$, where $\mathfrak{S}^{I}$ (respectively, $\mathfrak{S}^{II}$) only contains domains of Type I (respectively, Type II). Note that Type I and Type II are not mutually exclusive, but for those domains of both Type I and Type II, we can simply put them in $\mathfrak{S}^{II}$.
Before starting the proof, we summarize the dependency of some important constants that will be used in the proof as follows.
[…]
Here $A$ stands for the acylindrical constants. We draw an arrow from a constant $M$ to $N$ if $N$ depends on $M$. Remember that the dependency shown above is incomplete, but we hope it is helpful to the reader.
5.1. Quasi-trees from domains of Type I
In this subsection, we are going to prove the following proposition.
Proposition 5.1.
There exists a finite-index subgroup $H<G$ satisfying the following. For any sufficiently large constant $R$ and any $T>0$, there exist quasi-trees $\mathcal{T}_{1},\dots,\mathcal{T}_{n}$ such that $H$ acts on $\prod_{j=1}^{n}\mathcal{T}_{j}$ diagonally and for any choice of base points $o_{j}\in\mathcal{T}_{j}$
$\sum_{V\in\mathfrak{S}^{I}}\operatorname{\mathsf{d}}_{V}^{T,R}(1,h)\preceq\sum_{j=1}^{n}\operatorname{\mathsf{d}}_{\mathcal{T}_{j}}(o_{j},ho_{j})$
for any $h\in H$.
First, we recall the following lemma.
Lemma 5.2.
[BBF21, Prop. 3.4] Let a group $H$ act on a $\delta$–hyperbolic space $Y$. Assume that the image of $H$ in $\text{Isom}(Y)$ is acylindrical. Consider a loxodromic element $g\in H$ and the collection $\mathbb{A}$ of all $H$–translates of a fixed $(\lambda,c)$–quasi-axis of $g$. Then there exists a constant $\theta>0$ depending only on $\lambda,c,\delta$ and the acylindrical constants such that for any $\gamma\in\mathbb{A}$, the set
$\{h\in H\mid\operatorname{\mathsf{diam}}(\pi_{\gamma}(h\gamma))\geq\theta\}$
is a finite union of double $EC(g)$–cosets.
Corollary 5.3.
Let $U\in\mathfrak{S}^{I}$. Consider a $(\lambda,c)$–quasi-axis $\gamma\subset\mathcal{C}V$ for the acylindrical action of $G_{U}$. Then there exists $\theta>0$, only depending on $\lambda,c,E$ and the acylindrical constants, and a finite-index subgroup $G_{\gamma}<G$ such that every translate of $\gamma$ by an element of $G_{\gamma}\cap\operatorname{Stab}_{G}(V)$ either has finite Hausdorff distance with $\gamma$ or has $\theta$–bounded projection to $\gamma$.
Proof.
It is clear by Lemma 5.2 and separability of quasi-axes. ∎
By definition of relative HHGs, $\mathfrak{S}^{I}$ consists of finitely many $G$–orbits, so acylindrical constants for $U\in\mathfrak{S}^{I}$ can be chosen uniformly. Thus, Lemma 4.2 provides uniform constants $\lambda\geq 1,c\geq 0$ for every $U\in\mathfrak{S}^{I}$. This further gives a uniform constant $\theta>0$ by Corollary 5.3.
Let $\mathcal{U}$ be a $G$–representative set of $\mathfrak{S}^{I}$ such that $1\in\mathbf{P}_{U}$ for any $U\in\mathcal{U}$. Let $U\in\mathcal{U}$. Let $T>0$ and let $K>0$ be a sufficiently large constant that will be decided by Lemma 5.4. Lemma 4.2 provides a $G_{U}$–finite collection $\mathcal{A}_{U}=\mathcal{A}_{U}^{T,R}$ of $(\lambda,c)$–quasi-axes, where $R$ is provided by Proposition 4.4. By Corollary 5.3, we can find a finite-index subgroup $H_{U}<G$ such that for any $\gamma\in\mathcal{A}_{U}$ and $h\in H_{U}\cap\operatorname{Stab}_{G}(U)$, either $\operatorname{\mathsf{d}}_{Haus}(h\gamma,\gamma)<\infty$ or $\operatorname{\mathsf{diam}}\pi_{\gamma}(h\gamma)<\theta$. For any $g\in G$, we define $\mathcal{A}_{gU}:=\{g\gamma~{}|~{}\gamma\in\mathcal{A}_{U}\}$.
Let $\mathcal{A}:=\bigsqcup_{U\in\mathfrak{S}}\mathcal{A}_{U}$ and $H:=\bigcap_{U\in\mathcal{U}}H_{U}$. Since $\mathcal{U}$ is finite, $H$ is of finite index in $G$. By adding finitely many domains to $\mathcal{U}$ so that there is one representative for each $H$–orbit on $\mathfrak{S}^{I}$, we obtain an $H$–representative set $\tilde{\mathcal{U}}$ of $\mathfrak{S}^{I}$. We still assume that $1\in\mathbf{P}_{U}$ for any $U\in\tilde{\mathcal{U}}$. Let $\{\gamma_{1},\dots,\gamma_{n}\}$ be an $H$–representative set of $\mathcal{A}$. We assume that every representative $\gamma_{j}$ is contained in some $U\in\tilde{\mathcal{U}}$. Let $\mathcal{A}_{j}\subset\mathcal{A}$ be the $H$–orbit of $\gamma_{j}$.
Without loss of generality, we assume that $H$ is colorable instead of virtually colorable. Thus, the $H$–orbit of any domain is pairwise transverse. By [HP22, Lemma 3.4], every $H$–orbit of $\mathfrak{S}^{I}$ is an $H$–projection system with constant $s_{0}+4E$, where $s_{0}$ is the constant provided by Theorem 2.6. Thus, every $\mathcal{A}_{j}$ is an $H$–projection system with a uniform projection constant $\xi=\xi(\theta,s_{0},\lambda,c,E)$ by Theorem 2.4. The projections defined there will be denoted by $\Pi_{\gamma}$.
Using Theorem 2.1, we obtain modified projections $\Pi^{\prime}_{\gamma}$ such that $(\mathcal{A}_{j},\{\Pi^{\prime}_{\gamma}\})$ satisfies the strong projection axioms with constant $\xi^{\prime}=\xi^{\prime}(\xi)$ and that $\Pi_{\gamma}(\alpha)$ and $\Pi^{\prime}_{\gamma}(\alpha)$ are apart from each other within a uniform Hausdorff distance $\epsilon=\epsilon(\xi)$. For any $K^{\prime}\geq 4\xi^{\prime}$, $\mathcal{C}_{K^{\prime}}\mathcal{A}_{j}$ is a quasi-tree by Theorem 2.2. The following lemma is an estimate via the orbit map between the projections $\pi^{H}_{\gamma}$ in the relative HHG stucture and the projections $\Pi^{\prime}_{\gamma}$ in the quasi-tree $\mathcal{C}_{K^{\prime}}\mathcal{A}_{j}$.
Lemma 5.4.
Fix a base point $o_{j}\in\gamma_{j}$ for each $j=1,\dots,n$. There exists a sufficiently large constant $K^{\prime}=K^{\prime}(\xi,\lambda,c,E)$ and a constant $\Delta>0$ such that if $K\geq 2K^{\prime}$, then
$\sum_{\gamma\in\mathcal{A}_{j}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}^{H}_{\gamma}(1,h)\right\}\kern-3.01385pt\right\}_{K}\leq 8\operatorname{\mathsf{d}}_{\mathcal{C}_{K^{\prime}}\mathcal{A}_{j}}(o_{j},ho_{j})+\Delta$
for any $h\in H$ and any $j=1,\dots,n$.
For the proof of Lemma 5.4, we need the following lemma.
Lemma 5.5.
For any constants $A,B\geq 0$ and constants $L,M>0$,
$\frac{\left\{\kern-3.01385pt\left\{A+B\right\}\kern-3.01385pt\right\}_{L+M}}{L+M}\leq\frac{\left\{\kern-3.01385pt\left\{A\right\}\kern-3.01385pt\right\}_{L}}{L}+\frac{\left\{\kern-3.01385pt\left\{B\right\}\kern-3.01385pt\right\}_{M}}{M}.$
Proof.
Assume that $A+B\geq L+M$. First, if $A<L$ then $B>M$. Thus, $\left\{\kern-3.01385pt\left\{B\right\}\kern-3.01385pt\right\}_{M}=B\geq\frac{M}{L+M}(A+B)$. Second, if $B<M$ then $A>L$ and the same argument holds. Finally, if $A\geq L$ and $B\geq M$, then all thresholds are reached and the inequality is obviously true. ∎
Proof of Lemma 5.4.
For simplicity, we use $|p-q|$ to mean the distance between two points $p,q$ in the same space. Assume that $\gamma_{j}\subset\mathcal{C}U$ for some $U\in\tilde{\mathcal{U}}$. Since $1\in\mathbf{P}_{U}$, we have $|\pi_{gU}(1)-\rho^{U}_{gU}|\leq E$ for any $g\in H-\operatorname{Stab}_{H}(U)$ by definition of standard product regions.
Assume $\gamma=g\gamma_{j}$. By hyperbolicity, there exists a constant $F=F(E,\lambda,c)$ such that if $\gamma\subset\mathcal{C}gU\neq\mathcal{C}U$ then $|\pi^{H}_{\gamma}(1)-\Pi^{\prime}_{\gamma}(o_{j})|\leq|\pi_{gU}(1)-\rho^{U}_{gU}|+F+\epsilon\leq E+F+\epsilon$ Let $M>\xi^{\prime}+E+F+\epsilon$ and define $\delta_{\gamma}(h)=|\pi^{H}_{\gamma}(h)-\Pi^{\prime}_{\gamma}(ho_{j})|$. We see that if $\delta_{\gamma}(h)=\delta_{h^{-1}\gamma}(1)\geq M$ then $\gamma\subset\mathcal{C}hU$. Thus, for a fixed $h\in H$, there are only finitely many $\gamma\in\mathcal{A}_{j}$ such that $\delta_{\gamma}(h)\geq M$ by projection axiom (P2).
Let $K^{\prime}>2M+4\xi^{\prime}$. Define $D_{\gamma}(1,h)=|\Pi^{\prime}_{\gamma}(o_{j})-\Pi^{\prime}_{\gamma}(ho_{j})|$. By the triangle inequality and Lemma 5.5, we obtain that
$\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}^{H}_{\gamma}(1,h)\right\}\kern-3.01385pt\right\}_{K^{\prime}+2M}\leq\frac{K^{\prime}+2M}{K^{\prime}}\left\{\kern-3.01385pt\left\{D_{\gamma}(1,h)\right\}\kern-3.01385pt\right\}_{K^{\prime}}+\frac{K^{\prime}+2M}{2M}\left\{\kern-3.01385pt\left\{\delta_{\gamma}(1)+\delta_{\gamma}(h)\right\}\kern-3.01385pt\right\}_{2M}.$
Therefore,
$\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}^{H}_{\gamma}(1,h)\right\}\kern-3.01385pt\right\}_{K}\leq 2\left\{\kern-3.01385pt\left\{D_{\gamma}(1,h)\right\}\kern-3.01385pt\right\}_{K^{\prime}}+\frac{K^{\prime}+2M}{M}\left\{\kern-3.01385pt\left\{\delta_{\gamma}(1)\right\}\kern-3.01385pt\right\}_{M}.$
Summing over $\gamma\in\mathcal{A}_{j}$, we obtain that
$\sum_{\gamma\in\mathcal{A}_{j}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}^{H}_{\gamma}(1,h)\right\}\kern-3.01385pt\right\}_{K}$ $\leq 2\sum_{\gamma\in\mathcal{A}_{j}}\left\{\kern-3.01385pt\left\{D_{\gamma}(1,h)\right\}\kern-3.01385pt\right\}_{K^{\prime}}+\frac{K^{\prime}+2M}{M}\sum_{\gamma\in\mathcal{A}_{j}}\left\{\kern-3.01385pt\left\{\delta_{\gamma}(1)\right\}\kern-3.01385pt\right\}_{M}.$
By discussions above, $\Delta_{j}=\frac{K^{\prime}+2M}{M}\sum_{\gamma\in\mathcal{A}_{j}}\left\{\kern-3.01385pt\left\{\delta_{\gamma}(1)\right\}\kern-3.01385pt\right\}_{M}$ is a finite constant. Let $\Delta=\max_{1\leq j\leq n}\Delta_{j}$. We conclude by Theorem 2.3.
Now we finish the proof of Proposition 5.1.
Proof of Proposition 5.1.
Let $K$ and $K^{\prime}$ be the constant provided by Lemma 5.4. Let $T>0$. From discussions before Lemma 5.4, we know that the collection $\mathcal{A}$ of quasi-axes provided by Proposition 4.4 forms $n$ quasi-trees $\mathcal{C}_{K^{\prime}}\mathcal{A}_{j}$, $j=1,\dots,n$. Moreover, Proposition 4.4 tells us that there exists $R>0$ such that
$\sum_{V\in\mathfrak{S}^{I}}\operatorname{\mathsf{d}}_{V}^{T,R}(1,h)\preceq\sum_{\gamma\in\mathcal{A}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}^{H}_{\gamma}(1,h)\right\}\kern-3.01385pt\right\}_{K}.$
Finally, we conclude by Lemma 5.4. ∎
5.2. Quasi-trees from domains of Type II
In this subsection, we prove an analogue of Proposition 5.1 for domains of Type II.
Let $H<G$ be the subgroup provided by Proposition 5.1. Let $\mathcal{V}$ be an $H$–representative set of $\mathfrak{S}^{II}$ such that $1\in\mathbf{P}_{V}$ for any $V\in\mathcal{V}$. For $V\in\mathcal{V}$ we write $[V]$ to mean its $H$–orbits. By colorability and [HP22, Lemma 3.4], every $[V]$ is an $H$–projection system with constant $s_{0}+4E$.
Fix any $V\in\mathcal{V}$. By property (QT’) of the action $\operatorname{Stab}_{G}(V)\curvearrowright\mathcal{C}V$, there exist quasi-trees $T_{V}^{i}$ along with $\operatorname{Stab}_{H}(V)$–equivariant maps $\iota^{i}_{V}:\mathcal{C}V\to T_{V}^{i}$ for $i=1,\dots,n_{V}$ such that
$\prod_{i=1}^{n_{V}}\iota_{V}^{i}:\mathcal{C}V\to\prod_{i=1}^{n_{V}}T_{V}^{i}$
is a $(\lambda^{\prime},c^{\prime})$–quasi-isometric embedding. In particular, $\iota^{i}_{V}$ is $(\lambda^{\prime},c^{\prime})$–coarsely Lipschitz. Fix $i\in\{1,\dots,n_{V}\}$. It is conventional to extend the map $\iota^{i}_{V}$ on $[V]$ in an $H$–equivariant way. It means that we can construct a collection of quasi-trees $\mathbf{T}_{[V]}^{i}=\{T_{U}^{i}~{}|~{}U\in[V]\}$ with an $H$–action and a collection of coarsely Lipschitz maps $\{\iota_{U}^{i}:\mathcal{C}U\to T_{U}^{i}~{}|~{}U\in[V]\}$ such that the following diagram commutes for any $h\in H$ and $U\in[V]$.
[…]
Define $\Pi_{T_{hU}^{i}}(T_{U}^{i}):=\iota_{hU}^{i}(\rho^{U}_{hU})$ for any $U\in[V]$ and $hU\neq U$. Clearly, these projections are $H$–equivariant and the projection axioms pass to the collection $(\mathbf{T}_{[V]}^{i},\{\Pi_{T_{U}^{i}}\})$ under coarsely Lipschitz maps $\{\iota_{U}^{i}\}$. We modify the projections within an error $\epsilon$ such that $(\mathbf{T}_{[V]}^{i},\{\Pi^{\prime}_{T_{U}^{i}}\})$ satisfies the strong projection axioms with constant $\zeta=\zeta(s_{0},\lambda^{\prime},c^{\prime},E)$. For any $K^{\prime\prime}\geq 4\zeta$, $\mathcal{C}_{K^{\prime\prime}}\mathbf{T}_{[V]}^{i}$ is a quasi-tree by Theorem 2.2. Define $\operatorname{\mathsf{d}}_{T^{i}_{U}}(1,h):=|\iota^{i}_{U}(\pi_{U}(1))-\iota^{i}_{U}(\pi_{U}(h))|$ for any $U\in[V]$. For any $V\in\mathcal{V}$ and $i\in\{1,\dots,n_{V}\}$, fix a base point $o_{V}^{i}\in T_{V}^{i}$. The following proposition is an analogue of Proposition 5.1.
Proposition 5.6.
There exists a constant $K^{\prime\prime}=K^{\prime\prime}(\zeta,\lambda^{\prime},c^{\prime},E)$ such that if $R\geq 2K^{\prime\prime}$ then
$\sum_{U\in\mathfrak{S}^{II}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{U}(1,h)\right\}\kern-3.01385pt\right\}_{R}\preceq\sum_{V\in\mathcal{V}}\sum_{i=1}^{n_{V}}\operatorname{\mathsf{d}}_{\mathcal{C}_{K^{\prime\prime}}\mathbf{T}_{[V]}^{i}}(o_{V}^{i},ho_{V}^{i}).$
for any $h\in H$.
Proof.
Fix $V\in\mathcal{V}$ and $i\in\{1,\dots,n_{V}\}$. For any $U\in[V]-\{V\}$, we have $|\iota^{i}_{U}(\pi_{U}(1))-\Pi_{T^{i}_{U}}^{\prime}(o_{i})|\leq\lambda^{\prime}|\pi_{U}(1)-\rho^{V}_{U}|+c^{\prime}+\epsilon\leq\lambda^{\prime}E+c^{\prime}+\epsilon$. Let $K^{\prime\prime}>4\zeta+2(\lambda^{\prime}E+c^{\prime}+\epsilon)$. Similarly to the proof of Lemma 5.4, we estimate that
$\sum_{U\in[V]}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{T^{i}_{U}}(1,h)\right\}\kern-3.01385pt\right\}_{R}$$\preceq\left\{\kern-3.01385pt\left\{|\Pi_{T^{i}_{U}}^{\prime}(o_{i})-\Pi_{T^{i}_{U}}^{\prime}(ho_{i})|\right\}\kern-3.01385pt\right\}_{K^{\prime\prime}}$
$\preceq\operatorname{\mathsf{d}}_{\mathcal{C}_{K^{\prime\prime}}\mathbf{T}_{[V]}^{i}}(o_{V}^{i},ho_{V}^{i}).$
Here the first inequality follows from the triangle inequality and Lemma 5.5, and the second inequality is by Theorem 2.3. Since the map $\prod_{i=1}^{n_{U}}\iota_{U}^{i}:\mathcal{C}U\to\prod_{i=1}^{n_{U}}T_{U}^{i}$ is a quasi-isometric embedding for any $U\in\mathfrak{S}^{II}$, we conclude by summing the inequality over $1\leq i\leq n_{V}$ for all $V\in\mathcal{V}$. ∎
5.3. Proof of Theorem 1.3
Proof.
Let $R>0$ be sufficiently large that satisfies Proposition 5.1, Proposition 5.6 and Theorem 3.8. Let $T\geq R+2D_{0}$. By Proposition 5.1 and Proposition 5.6, there exists quasi-trees $\mathcal{T}_{1},\dots,\mathcal{T}_{m}$ such that $H$ acts on $\prod_{k=1}^{m}\mathcal{T}_{k}$ diagonally and for any choice of base points $o_{k}\in\mathcal{T}_{k}$ and any $h\in H$,
$\sum_{V\in\mathfrak{S}^{I}}\operatorname{\mathsf{d}}_{V}^{T,R}(1,h)+\sum_{U\in\mathfrak{S}^{II}}\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{U}(1,h)\right\}\kern-3.01385pt\right\}_{R}\preceq\sum_{k=1}^{m}\operatorname{\mathsf{d}}_{\mathcal{T}_{k}}(o_{k},ho_{k}).$
By definition of thick distance, $\operatorname{\mathsf{d}}_{U}^{T,R}(1,h)\leq\left\{\kern-3.01385pt\left\{\operatorname{\mathsf{d}}_{U}(1,h)\right\}\kern-3.01385pt\right\}_{R}$. Thus, by Theorem 3.8,
$\sum_{V\in\mathfrak{S}}\operatorname{\mathsf{d}}_{V}^{T,R}(1,h)\preceq\sum_{k=1}^{m}\operatorname{\mathsf{d}}_{\mathcal{T}_{k}}(o_{k},ho_{k})$
for any $h\in H$.
On the other hand, the orbit map from $H$ to $\prod_{k=1}^{m}\mathcal{T}_{k}$ is coarsely Lipschitz since $H$ is finitely generated. Therefore, $H$ embeds quasi-isometrically into $\prod_{k=1}^{m}\mathcal{T}_{k}$, which means that $H$ has property (QT’). Finally, we know that $G$ has property (QT) since property (QT) passes to any finite-index supergroup.
For the “moreover” part, first note that $G$ is coarse median for the same reason as [BHS19, Theorem 7.3]. The rest of the proof is just a combination of [HP22, Pet21]. The proof of [HP22, Proposition 3.9] can be naturally generalized to quasi-trees from domains of Type II. The proof in [Pet21, §3] for mapping class groups can be repeated verbatim to deal with quasi-trees from domains of Type I. ∎
For relative HHGs with only Type II domains, we obtain the following stronger theorem.
Theorem 5.7.
Let $(G,\mathfrak{S})$ be a relative HHG that is colorable. If every $U\in\mathfrak{S}$ is of Type II, then $G$ has property (QT’).
Proof.
The proof is just a simplified version of that of Theorem 1.3. Note that $G$ is colorable and that every domain is of Type II, so the finite-index subgroup $H$ in the above proof can be replaced with $G$ itself. ∎
6. A criterion for having separable quasi-axes
In this section, we provide a criterion for a relative HHG to have separable quasi-axes that is easy to use in application.
Definition 6.1 (Amalgamated direct product).
[NN50] A group $G$ is said to be the direct product of its subgroups $H$ and $K$ with an amalgamated subgroup $Z$ if (i) $G$ is generated by $H\cup K$, (ii) $H\cap K=Z$, (iii) $[H,K]=\{1\}$. It is denoted by $G=H\times_{Z}K$.
For an acylindrical action on a hyperbolic space, we have seen in Section 2.1 that every elementary closure is a virtual centralizer. The following lemma is inspired by the discussion in [BBF21, §4.3].
Lemma 6.2.
Let $G=H\times_{Z}K$ be an amalgamated direct product of groups. Assume that $G$ acts on a $\delta$–hyperbolic space $X$ with an acylindrical image $\bar{G}<\text{Isom}(X)$ such that $K$ is the kernel of this action and $H$ surjects onto $\bar{G}$ under this action. Then for any loxodromic element $h\in H$, the elementary closure $EC_{G}(h)$ contains a subgroup of index at most $2$ that is a centralizer of $h^{k}$ for some $k>0$. In particular, for any loxodromic element $g\in G$, the elementary closure $EC_{G}(g)$ is a virtual centralizer in $G$.
Proof.
Let $\theta:G\to\bar{G}$ be the quotient map. For any $g\in G$, we write $\bar{g}$ to mean $\theta(g)\in\bar{G}$. Let $h\in H$ be a loxodromic element. By definition, $EC_{G}(h)=\theta^{-1}(EC_{\bar{G}}(\bar{h}))$. Let $EC_{\bar{G}}^{+}(\bar{h})$ be the subgroup of $EC_{\bar{G}}(\bar{h})$ that fixes the limit points of the $\langle\bar{h}\rangle$–action on $\partial X$. Clearly, $EC_{\bar{G}}^{+}(\bar{h})$ has index at most $2$ in $EC_{\bar{G}}(\bar{h})$. In the same way, we define $EC_{G}^{+}(h)$, which has index at most $2$ in $EC_{G}(h)$. We only need to prove that $EC_{G}^{+}(h)$ is a centralizer of $h^{k}$ for some $k>0$.
Without loss of generality, assume that $h$ is primitive. Since $\bar{G}$ acts acylindrically, $EC_{\bar{G}}^{+}(\bar{h})$ is virtually cyclic and acts without flips. Thus, $EC_{\bar{G}}^{+}(\bar{h})$ is a finite extension of $\mathbb{Z}$ that fits into a short exact sequence
$1\to F\to EC_{\bar{G}}^{+}(\bar{h})\to\mathbb{Z}\to 1,$
where $F$ is finite and the subgroup $\langle\bar{h}\rangle$ surjects onto $\mathbb{Z}$. As $\langle\bar{h}\rangle$ acts on $F$ by conjugation, there exists $n>0$ such that the action of $\bar{h}^{n}$ is the identity, which means that $\bar{h}^{n}$ commutes with every element of $F$.
Consider the quotient map $\theta|_{EC_{G}^{+}(h)}:EC_{G}^{+}(h)\to EC_{\bar{G}}^{+}(\bar{h})$. Every element in $EC_{G}^{+}(h)$ can be written as a product of an element in $\theta^{-1}(F)$ and a power of $h$. Since $\bar{G}=\theta(H)$, we can further decompose $EC_{G}^{+}(h)$ into
(2) $EC_{G}^{+}(h)=K\cdot(H\cap\theta^{-1}(F))\cdot\langle h\rangle.$
Let $m=|F|$. For any $f\in H\cap\theta^{-1}(F)$, we have $[\bar{f},\bar{h}^{mn}]=1$. Thus, $[f,h^{mn}]\in K$. Since $f,h\in H$, a straightforward calculation shows that $[f,h^{mn}]=[f,h^{n}]^{m}=[f^{m},h^{n}]=1$. The last equation holds because $\bar{f}^{m}=1$, which implies $f^{m}\in K$. Therefore, $h^{mn}$ commutes with every element in $EC_{G}^{+}(h)$ by Equation (2). This implies $EC_{G}^{+}(h)\subset C_{H}(h^{mn})$.
On the other hand, any element in $H$ that commutes with $h^{mn}$ has an image in $\bar{H}$ that commutes with $\bar{h}^{mn}$. Thus, $C_{H}(h^{mn})\subset EC_{G}^{+}(h)$. In conclusion, we obtain $EC_{G}^{+}(h)=C_{H}(h^{mn})$. ∎
Definition 6.3.
For a relative HHG $(G,\mathfrak{S})$, we say a domain $V\in\mathfrak{S}$ has neat kernel if $\operatorname{Stab}_{G}(V)$ decomposes as an amalgamated direct product $H_{V}\times_{Z_{V}}K_{V}$ such that $K_{V}=\operatorname{Ker}(\theta_{V})$ and $\theta_{V}(H_{V})=G_{V}$.
Corollary 6.4.
Let $(G,\mathfrak{S})$ be a relative HHG that is residually finite. Let $V\in\mathfrak{S}$. If $V$ has hyperbolicity, acylindrical image and neat kernel, then $V$ has separable quasi-axes.
Proof.
Let $g\in\operatorname{Stab}_{G}(V)$ be a loxodromic element. Since $V$ has neat kernel, we can assume that $g\in H_{V}$ with $EC(g)$ not changed. By Lemma 6.2, $EC(g)$ is a virtual centralizer of $g^{k}$ for some $k>0$ in $\operatorname{Stab}_{G}(V)$. Every element that commutes with $g^{k}$ stabilizes $V$. Therefore, $EC(g)$ is a virtual centralizer of $g^{k}$ in $G$. It is known that a centralizer in a residually finite group is separable (see [BBF21, Lemma 2.1] or the proof of [Lon87, Propostion]). It is also known that a finite-index supergroup of a separable subgroup is still separable (easy to see from the profinite topology). Therefore, $V$ has separable quasi-axes. ∎
Corollary 6.5.
Let $(G,\mathfrak{S})$ be a relative HHG that is residually finite. Let $S\in\mathfrak{S}$ be the unique maximal domain. Then $S$ is of Type I.
Proof.
By [BHS17b, Theorem 14.3], $G$ acts on $\mathcal{C}S$ acylindrically. Now $\operatorname{Ker}(\theta_{V})$ is trivial so $S$ has neat kernel. Thus, $S$ has separable quasi-axes by Corollary 6.4. Moreover, $S$ has cobounded nested region because $G$ acts on $\mathcal{X}$ coboundedly. In conclusion, $S$ is of Type I. ∎
7. Applications
7.1. Mapping class groups
In this subsection, we explain how Theorem 1.3 applies to mapping class groups to recover the following theorem.
Theorem 7.1.
[BBF21, Theorem 1.2] Mapping class groups have property (QT).
Proof.
Let $\Sigma$ be a closed oriented surface with finitely many marked points, and $\mathcal{M}(\Sigma)$ be its marking complex [MM00]. Let $\mathfrak{S}$ be the collection of isotopy classes of closed essential non-pants subsurfaces of $\Sigma$, where disconnected subsurfaces are also allowed. Given any $V\in\mathfrak{S}$, let $\mathcal{C}V$ be the curve graph of $V$. (The curve graph of a disconnected surface is the join of the curve graphs of its components, and thus is bounded.) Then it is known that $(\mathcal{M}(\Sigma),\mathfrak{S})$ is an HHS. Moreover, the mapping class group $\operatorname{MCG}(\Sigma)$ is an HHG by virtue of its action on $(\mathcal{M}(\Sigma),\mathfrak{S})$ [BHS19, §11].
For any $V\in\mathfrak{S}^{\circ}$, let $V^{\perp}$ be its complementary (closed) surface in $\Sigma$ and let $\operatorname{MCG}(\Sigma;V)$ be the stabilizer of $V$ in $\operatorname{MCG}(\Sigma)$. If $V$ is connected and non-annular, let $\bar{V}$ be the surface obtained from $V$ by collapsing the components of $\partial V$ to marked points. By [FM12, Theorem 3.18] $\operatorname{MCG}(V)$ and $\operatorname{MCG}(V^{\perp})$ are subgroups of $\operatorname{MCG}(\Sigma;V)$. By [FM12, Proposition 3.20], $\operatorname{MCG}(V^{\perp})$ is exactly the kernel of the restriction homomorphism $\theta_{V}:\operatorname{MCG}(\Sigma;V)\to\operatorname{MCG}(\bar{V})$. Moreover, it is easy to see that $\operatorname{MCG}(V)$ commutes with $\operatorname{MCG}(V^{\perp})$. Thus, $\operatorname{MCG}(\Sigma;V)$ decomposes as an amalgamated direct product of $\operatorname{MCG}(V)$ and $\operatorname{MCG}(V^{\perp})$. This implies that $V$ has neat kernel.
It is known that $\operatorname{MCG}(\Sigma)$ is virtually colorable [BBF15, §5] and residually finite [Gro74]. We only need to prove that any $V\in\mathfrak{S}$ with $\mathcal{C}V$ not being a quasi-tree is of Type I. In this case, $V$ is connected and non-annular so $\mathbf{F}_{V}$ is coarsely $\mathcal{M}(V)$. Since $\operatorname{MCG}(\bar{V})$ acts coboundedly on $\mathcal{M}(V)$ and acts acylindrically on $\mathcal{C}V$ [Bow08], $V$ has cobounded nested region and acylindrical image. Furthermore, $V$ has separable quasi-axes by Corollary 6.4. Therefore, $V$ is of Type I.
In conclusion, mapping class groups have property (QT) by Theorem 1.3. ∎
7.2. Admissible graphs of groups
Admissible groups are introduced by Croke-Kleiner in [CK02], which generalize the fundamental groups of non-geometric $3$–dimensional graph manifolds.
Definition 7.2.
Let $\mathcal{G}=(\Gamma,\{G_{v}\},\{G_{e}\})$ be a graph of groups. We say $\mathcal{G}$ is admissible if the following hold:
  1. (1)
    $\Gamma$ is a finite graph with at least one edge.
  2. (2)
    Each vertex group ${G}_{v}$ has center $Z_{v}\cong\mathbb{Z}$, and ${H}_{v}:={G}_{v}/Z_{v}$ is a non-elementary hyperbolic group.
  3. (3)
    Every edge group ${G}_{e}$ is isomorphic to $\mathbb{Z}^{2}$.
  4. (4)
    If $e$ is an edge with $v=e^{+}$ and $w=e^{-}$, and $\tau_{e},\tau_{\bar{e}}$ are the edge monomorphisms, then the subgroup $\langle\tau_{e}^{-1}(Z_{v}),\tau_{\bar{e}}^{-1}(Z_{w})\rangle$ has finite index in ${G}_{e}$.
  5. (5)
    Let $e_{1}$ and $e_{2}$ be distinct edges entering a vertex $v$, and let $K_{i}\subset{G}_{v}$ be the image of the edge homomorphism $\tau_{e_{i}}$ for $i=1,2$. Then
    • for every $g\in G_{v}$, $gK_{1}g^{-1}$ is not commensurable with $K_{2}$;
    • for every $g\in G_{v}-K_{i}$, $gK_{i}g^{-1}$ is not commensurable with $K_{i}$.
A group $G$ is admissible if it is the fundamental group of an admissible graph of groups.
Every admissible group $G$ has a (combinatorial) HHG structure by [HRSS24, Theorem 1.4]. According to the classification of simplices by [HRSS24, Lemma 6.2], if $\Delta\perp g\Delta$, where $\Delta$ corresponds to an unbounded hyperbolic space, then $\Delta$ is of Type $8$ and $g$ exchanges two adjacent vertices in the Bass-Serre tree (see [BHMS20, Definition 1.11] for definition of orthogonality in a combinatorial HHS). Therefore, it is easy to see that $G$ has a subgroup of index at most $2$ that is colorable (see [NY23, Lemma 4.6] for example). This shows the virtual colorability of $G$. In particular, every non-geometric graph manifold group has a virtually colorable HHG structure with all associated hyperbolic spaces being quasi-trees. Thus, non-geometric graph manifold groups have property (QT) by [HP22, Theorem 3.1] or Theorem 1.3. However, in the HHG structure of an admissible group, associated hyperbolic spaces are not necessarily quasi-trees. As an application of Theorem 1.3, we show that property (QT) still holds true in this case if we assume $G$ to be residually finite.
Theorem 7.3.
Let $\mathcal{G}=(\Gamma,\{G_{v}\},\{G_{e}\})$ be an admissible graph of groups, and let $G=\pi_{1}\mathcal{G}$. If $G$ is residually finite, then $G$ has property (QT).
Proof.
According to the classification of simplices by [HRSS24, Lemma 6.2], any simplex that is not of Type $7$ corresponds to a quasi-tree so it is a domain of Type II. Thus, we only need to check that simplices of Type $7$ are of Type I. The stabilizer of such a simplex $\Delta$ is exactly a vertex group $G_{v}$ that acts on $\mathcal{C}(\Delta)$ with image $H_{v}=G_{v}/Z_{v}$. Now $\mathcal{C}(\Delta)$ is coarsely the hyperbolic space obtained by coning off $H_{v}$ as a relatively hyperbolic group, and $\mathbf{F}_{\Delta}$ is coarsely $H_{v}$ itself. Therefore, acylindrical image and cobounded nested region hold true (see [Osi16, Proposition 5.2] for acylindricity). Since $G_{v}$ is a central extension of $H_{v}$ by $Z_{v}$, $\Delta$ has neat kernel. Therefore, we conclude by Corollary 6.4 and Theorem 1.3. ∎
There is another approach to property (QT) of non-geometric graph manifold groups in [HNY21]. For graph manifolds with nonempty boundary, they actually prove in a more general setting. We say a CKA group is an admissible group that admits a geometric action on a complete proper CAT(0) space. As a corollary of Theorem 7.3, we recover the following theorem.
Corollary 7.4.
[HNY21, Theorem 1.3] Let $G$ be a CKA group where for every vertex $v$ the central extension $1\to Z_{v}\to G_{v}\to H_{v}\to 1$ has an omnipotent hyperbolic quotient group $H_{v}$. Then $G$ has property (QT).
For definition of omnipotence, we refer the reader to [Wis00]. Note that if every hyperbolic group is residually finite, then every hyperbolic group is omnipotent by [Wis00, Remark 3.4]. Under the assumption of Corollary 7.4, the central extension associated with any vertex virtually splits by [BH99, Theorem II.7.1]. Therefore, Theorem 7.3 implies Corollary 7.4 due to the following lemma.
Lemma 7.5.
Let $G$ be an admissible group where for every vertex $v$ the central extension $1\to Z_{v}\to G_{v}\to H_{v}\to 1$ virtually splits and the hyperbolic quotient group $H_{v}$ is omnipotent. Then $G$ is residually finite.
We omit the proof of Lemma 7.5 since it is almost the same as the proof of residual finiteness for graph manifold groups by Hempel [Hem87]. The reader can also see [Ngu24] for an improved result.
7.3. Hyperbolic–$2$–decomposable groups
We say a group $G$ is hyperbolic–$2$–decomposable if $G$ splits as a graph of hyperbolic groups with $2$-ended edge groups.
Theorem 7.6.
Let $G$ be a residually finite hyperbolic–$2$–decomposable group. Then $G$ has property (QT) if and only if $G$ does not contain any distorted element.
Proof.
If $G$ has property (QT), then $G$ does not contain any distorted element by [HNY21, Lemma 2.5]. Now assume that $G$ does not contain any distorted element. Let $\mathfrak{S}$ be the HHG structure of $G$ given by [RS20]. By construction, there is no orthogonality in $\mathfrak{S}$. Thus, $G$ is colorable. Let $U\in\mathfrak{S}$. Then $\mathcal{C}U$ is either a quasi-tree or a hyperbolic space obtained by coning off a vertex group $G_{v}$ as a relatively hyperbolic group. In the former case, $U$ is of Type II. Now we only need to consider the latter case. Similarly to Theorem 7.3, we have $\operatorname{Stab}_{G}(U)=G_{v}$ and $\mathbf{F}_{U}$ is coarsely $G_{v}$ itself. It is easy to see that acylindrical image, cobounded nested region and neat kernel hold true. Therefore, we conclude by Corollary 6.4 and Theorem 1.3. ∎
Similarly to Lemma 7.5, if $G$ is a hyperbolic–$2$–decomposable group without any distorted element such that every vertex group is omnipotent, then $G$ is residually finite (see [Wis00, §4]).
7.4. HHGs with all associated hyperbolic spaces being quasi-trees except the maximal one
Let $G$ be either of the following groups.
  • An Artin group of large and hyperbolic type;
  • The $\pi_{1}(\Sigma)$–extension group of a lattice Veech group in the mapping class group $\operatorname{MCG}(\Sigma)$ of a closed surface $\Sigma$.
As shown in [HMS24] and [DDLS24] respectively, $G$ is a virtually colorable HHG. Moreover, the associated hyperbolic spaces of $G$ are all quasi-trees except the maximal one. By Corollary 6.5 and Theorem 1.3, $G$ has property (QT) if $G$ is residually finite. We ask
Question 7.7.
When is $G$ residually finite?
It is proved in [Jan22] that any $3$–generator Artin groups with labels $\geq 4$ except for $(2m+1,4,4)$ for any $m\geq 2$ is residually finite. As a corollary, we obtain the following.
Theorem 7.8.
Any $3$–generator Artin group with labels $\geq 4$ except for $(2m+1,4,4)$ for any $m\geq 2$ has property (QT).
7.5. Graph products
Definition 7.9 (Graph product).
Let $\Gamma$ be a finite simplicial graph with the vertex set $V(\Gamma)$ and the edge set $E(\Gamma)$. Each vertex $v\in V(\Gamma)$ is labeled by a group $G_{v}$. The graph product $G_{\Gamma}$ is the group
[…]
We call $G_{v}$ the vertex groups of the graph product $G_{\Gamma}$.
Theorem 7.10.
Let $G_{\Gamma}$ be a graph product of groups with every vertex group having property (QT’). Then $G_{\Gamma}$ has property (QT’).
Proof.
[…]